### v4

# Arithmetic Properties of Several Generalized-Constant Sequences, with Implications for $\boldsymbol{\Gamma^{\left(n\right)}\left(1\right)}$

Abstract.
Neither the Euler-Mascheroni constant, $\gamma=0.577215\ldots$, nor the Euler-Gompertz constant, $\delta=0.596347\ldots$, is currently known to be irrational. However, it has been proved that at least one of them is transcendental. The two constants are related through a well-known equation of Hardy, equivalent to $\gamma+\delta/e=\textrm{Ein}\left(1\right)$, which recently has been generalized to $\gamma^{\left(n\right)}+\delta^{\left(n\right)}/e=\eta^{\left(n\right)},\>n\geq 0$ for sequences of constants $\gamma^{\left(n\right)}$, $\delta^{\left(n\right)}$, and $\eta^{\left(n\right)}$ (given respectively by raw, conditional, and partial moments of the Gumbel(0,1) probability distribution). Investigating the $\gamma^{\left(n\right)}$ through recurrence relations (where $\gamma^{\left(0\right)}=1$ and $\gamma^{\left(1\right)}=\gamma$), we find that at least one of the pair $\{\gamma,\gamma^{\left(2\right)}\}$ and – conditional on a realistic conjecture verified for $2\leq n\leq 26$ – at least two of each set $\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\}$ are transcendental, implying that the $\gamma^{\left(n\right)}$ are transcendental infinitely often (with analogous results for the sequence $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$). We then show, via a theorem of Shidlovskii, that the $\eta^{\left(n\right)}$ are algebraically independent, and therefore transcendental, for all $n\geq 0$, implying that at least one of each pair, $\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e\}$ and $\{\gamma^{\left(n\right)},\delta^{\left(n\right)}\}$, and at least two of the triple $\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e,\delta^{\left(n\right)}\}$, are transcendental for all $n\geq 1$. Further analysis of the $\gamma^{\left(n\right)}$ and $\eta^{\left(n\right)}$ reveals that the values $\delta^{\left(n\right)}/e$ are transcendental infinitely often with positive asymptotic density. Finally, we provide parallel results for the sequences $\widetilde{\delta}^{\left(n\right)}$ and $\widetilde{\eta}^{\left(n\right)}$ satisfying the “non-alternating analogue” equation $\gamma^{\left(n\right)}+\widetilde{\delta}^{\left(n\right)}/e=\widetilde{\eta}^{\left(n\right)}$.
Key words and phrases:
Euler-Mascheroni constant; Euler-Gompertz constant; Gumbel distribution; irrationality; transcendence; algebraic independence; gamma-function derivatives.
2000 Mathematics Subject Classification:
11J81, 60E05
1. Introduction
1.1. Three Fundamental Sequences
Let $\gamma=0.577215\ldots$ denote the Euler-Mascheroni constant and $\delta=0.596347\ldots$ denote the Euler-Gompertz constant. Although neither $\gamma$ nor $\delta$ has been shown to be irrational, Aptekarev [1] was the first to note their disjunctive irrationality; that is, at least one of the two numbers must be irrational. [^1] Aptekarev’s observation in [1] was based on earlier work of Shidlovskii [2] and Mahler [3] that implied, but did not explicitly state, the disjunctive irrationality of $\left\{\gamma,\delta\right\}$ . More recently, Rivoal [4] strengthened this result to show the disjunctive transcendence of $\left\{\gamma,\delta\right\}$. [^2] See Lagarias [5] for a comprehensive treatment of results involving these two constants.
The two constants are related by the intriguing equation
$\delta=-e\left[\gamma-{\displaystyle\sum_{k=1}^{\infty}\dfrac{\left(-1\right)^{k+1}}{k\cdot k!}}\right],$ (1)
introduced by Hardy [6], which can be rewritten as
$\gamma+\dfrac{\delta}{e}=-{\displaystyle\sum_{k=1}^{\infty}\dfrac{\left(-1\right)^{k}}{k\cdot k!}}=\textrm{Ein}\left(1\right),$ (2)
where $\textrm{Ein}\left(z\right)=-{\textstyle\sum_{k=1}^{\infty}\left(-z\right)^{k}/\left(k\cdot k!\right)}$ is the entire exponential-integral function. [^3] Both (1) and (2) follow immediately from setting z=1z=-1 in the power-series expansion of the exponential-integral function (using the principal branch for $z\in\mathbb{R}_{<0}$), $\textrm{Ei}\left(z\right)=\gamma+\ln\left|z\right|+{\textstyle\sum_{k=1}^{\infty}}z^{k}/\left(k\cdot k!\right)$, and recognizing that $\delta=-e\textrm{Ei}\left(-1\right)$. Although it is known that $\gamma+\delta/e$ is transcendental, this fact is infrequently mentioned in the literature. [^4] Aptekarev [1] and Lagarias [5] (Section 3.16, p. 608) attributed the observation that 11/e1-1/e and $\left(\gamma+\delta/e\right)$ are algebraically independent, and therefore transcendental, to a more general result asserted (without proof) by Mahler [3]. (However, Aptekarev [1] contains a typographical error, referring to $\left(\gamma-\delta/e\right)$ instead of $\left(\gamma+\delta/e\right)$.)
Recently, we provided the following probabilistic interpretation of (2): [^5] See Powers [7].
$\textrm{E}_{X}\left[X\right]-\textrm{E}_{X}\left[X\mid X\leq 0\right]\Pr\left\{X\leq 0\right\}=\textrm{E}_{X}\left[X^{+}\right],$ (3)
where $X\sim\textrm{Gumbel}\left(0,1\right)$ [^6] The cumulative distribution function of $\textrm{Gumbel}\left(0,1\right)$ is given by $F_{X}\left(x\right)=\exp\left(-e^{-x}\right),\>x\in\mathbb{R}$, with mean $\gamma$ and variance $\pi^{2}/6$. and $X^{+}=\max\left\{X,0\right\}$, so that
$\gamma=\textrm{E}_{X}\left[X\right]={\displaystyle\int_{-\infty}^{\infty}}x\exp\left(-x-e^{-x}\right)dx,$
$\delta=-\textrm{E}_{X}\left[X\mid X\leq 0\right]=-e{\displaystyle\int_{-\infty}^{0}}x\exp\left(-x-e^{-x}\right)dx,$
$\dfrac{1}{e}=\Pr\left\{X\leq 0\right\}={\displaystyle\int_{-\infty}^{0}}\exp\left(-x-e^{-x}\right)dx,$
and
$\sum_{k=1}^{\infty}\dfrac{\left(-1\right)^{k}}{k\cdot k!}=\textrm{E}_{X}\left[X^{+}\right]={\displaystyle\int_{0}^{\infty}}x\exp\left(-x-e^{-x}\right)dx.$
The identity in (3) suggests the natural generalization
$\textrm{E}_{X}\left[X^{n}\right]-\textrm{E}_{X}\left[X^{n}\mid X\leq 0\right]\Pr\left\{X\leq 0\right\}=\textrm{E}_{X}\left[\left(X^{+}\right)^{n}\right],\quad n\in\mathbb{Z}_{\geq 0}$ (4)
(with $\left(X^{+}\right)^{0}\coloneqq\mathbf{1}_{\left\{X>0\right\}}$), which can be expressed in the manner of (2) as
$\gamma^{\left(n\right)}+\dfrac{\delta^{\left(n\right)}}{e}=-n!{\displaystyle\sum_{k=1}^{\infty}\dfrac{\left(-1\right)^{k}}{k^{n}\cdot k!}},$ (5)
where
$\gamma^{\left(n\right)}=\textrm{E}_{X}\left[X^{n}\right]={\displaystyle\int_{-\infty}^{\infty}}x^{n}\exp\left(-x-e^{-x}\right)dx,$
$\delta^{\left(n\right)}=-\textrm{E}_{X}\left[X^{n}\mid X\leq 0\right]=-e{\displaystyle\int_{-\infty}^{0}}x^{n}\exp\left(-x-e^{-x}\right)dx,$ (6)
and
$n!{\displaystyle\sum_{k=1}^{\infty}\dfrac{\left(-1\right)^{k}}{k^{n}\cdot k!}}=\textrm{E}_{X}\left[\left(X^{+}\right)^{n}\right]={\displaystyle\int_{0}^{\infty}}x^{n}\exp\left(-x-e^{-x}\right)dx.$
For ease of exposition, we will set
$\eta^{\left(n\right)}=-n!{\displaystyle\sum_{k=1}^{\infty}\dfrac{\left(-1\right)^{k}}{k^{n}\cdot k!}}$ (7)
so that (5) may be rewritten as
$\gamma^{\left(n\right)}+\dfrac{\delta^{\left(n\right)}}{e}=\eta^{\left(n\right)}.$ (8)
Further, we will refer to the sequences $\gamma^{\left(n\right)}$, $\delta^{\left(n\right)}$, and $\eta^{\left(n\right)}$ as the generalized Euler-Mascheroni, Euler-Gompertz, and Eta constants, respectively (with $\gamma^{\left(1\right)}=\gamma$, $\delta^{\left(1\right)}=\delta=-e\textrm{Ei}\left(-1\right)$, and $\eta^{\left(1\right)}=\eta=\textrm{Ein}\left(1\right)$ denoting the ordinary Euler-Mascheroni, Euler-Gompertz, and Eta constants). [^7] Note that the generalized Euler-Mascheroni constants, $\gamma^{\left(n\right)}$, are indexed via superscripts to distinguish them from the well-known sequence of Stieltjes constants, commonly denoted by $\gamma_{n}$. Furthermore, the generalized Eta constants, $\eta^{\left(n\right)}$, should not be confused with the eta functions of Dirichlet, Dedekind, or Weierstrass.
1.2. Asymptotic Behavior
Table 1 presents the first 16 values of the generalized Euler-Mascheroni, Euler-Gompertz, and Eta constants. As is clear from a brief inspection, both the $\gamma^{\left(n\right)}$ and $\eta^{\left(n\right)}$ increase approximately factorially, whereas the $\delta^{\left(n\right)}$, which alternate in sign, grow super-exponentially but sub-factorially in magnitude.
The following proposition quantifies the asymptotic behavior underlying these observations.
Proposition 1.
As $n\rightarrow\infty$:
(i) $\gamma^{\left(n\right)}=n!\left(1-1/2^{n+1}+O\left(1/3^{n}\right)\right)$;
(ii) $\left|\delta^{\left(n\right)}\right|=e\left[W\left(n\right)\right]^{n}\exp\left(-n/W\left(n\right)\right)\sqrt{2\pi n/\left(W\left(n\right)+1\right)}\left(1+o\left(1\right)\right)$, where $W\left(n\right)$ denotes the principal branch of the Lambert W function for $n>0$; and
(iii) $\eta^{\left(n\right)}=n!\left(1-1/2^{n+1}+O\left(1/3^{n}\right)\right)$.
Proof.
See the Appendix. ∎
Table 1. Values of $\gamma^{\left(n\right)}$, $\delta^{\left(n\right)}$, and $\eta^{\left(n\right)}$ for $n\in\left\{0,1,\ldots,15\right\}$
$\boldsymbol{n}$ $\boldsymbol{\gamma^{\left(n\right)}}$ $\boldsymbol{\delta^{\left(n\right)}}$ $\boldsymbol{\eta^{\left(n\right)}}$
0 $1.0$ 1.0-1.0 $0.6321205588\ldots$
11 $0.5772156649\ldots$ $0.5963473623\ldots$ $0.7965995992\ldots$
22 $1.9781119906\ldots$ $0.5319307700\ldots$ $1.7824255962\ldots$
33 $5.4448744564\ldots$ $0.5806819508\ldots$ $5.6584954080\ldots$
44 $23.5614740840\ldots$ $0.7222515339\ldots$ $23.2957725933\ldots$
55 $117.8394082683\ldots$ $0.9875880596\ldots$ $118.2027216118\ldots$
66 $715.0673625273\ldots$ $1.4535032853\ldots$ $714.5326485509\ldots$
77 $5,019.8488726298\ldots$ $2.2708839827\ldots$ $5,020.6842841603\ldots$
88 $40,243.6215733357\ldots$ $3.7298791058\ldots$ $40,242.2494274946\ldots$
99 $362,526.2891146549\ldots$ $6.3945118625\ldots$ $362,528.6415241055\ldots$
$10$ $\qquad 3,627,042.4127568947\ldots\qquad$ $11.3803468877\ldots$ $\qquad 3,627,038.2261612415\ldots\qquad$
$11$ $39,907,084.1514313358\ldots$ $20.9346984188\ldots$ $39,907,091.8528764918\ldots$
$12$ $478,943,291.7651829432\ldots$ $39.6671864816\ldots$ $478,943,277.1724405288\ldots$
$13$ $6,226,641,351.5460642549\ldots$ $77.1984745660\ldots$ $6,226,641,379.9457959376\ldots$
$14$ $87,175,633,810.7084156319\ldots$ $153.9437943882\ldots$ $87,175,633,754.0756585806\ldots$
$15$ $1,307,654,429,495.7941762096\ldots$ $313.9164765016\ldots$ $1,307,654,429,611.2775941595\ldots$
1.3. Outline of Results
In Section 2, we begin our study of the arithmetic properties of the three generalized-constant sequences by investigating recurrence relations associated with the $\gamma^{\left(n\right)}$. These recurrences reveal that at least one element of the pair $\left\{\gamma,\gamma^{\left(2\right)}\right\}$ and – conditional on a realistic conjecture computationally verified for $2\leq n\leq 26$ – at least two elements of each set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$ are transcendental, implying that the $\gamma^{\left(n\right)}$ are transcendental infinitely often (with analogous results for the sequence $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$). Next, in Section 3, we apply a theorem of Shidlovskii (1989) to demonstrate the algebraic independence, and therefore transcendence, of the $\eta^{\left(n\right)}$ for all $n\geq 0$. This implies that at least one element of each pair, $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e\right\}$ and $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}\right\}$, and at least two elements of the triple $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e,\delta^{\left(n\right)}\right\}$, are transcendental for all $n\geq 1$. Moreover, further analysis of the $\gamma^{\left(n\right)}$ and $\eta^{\left(n\right)}$ shows that the values $\delta^{\left(n\right)}/e$ are transcendental infinitely often, with the density of the set of transcendental terms having asymptotic lower bound $1/2$. Finally, in Section 4 we consider two additional sequences, $\widetilde{\delta}^{\left(n\right)}$ and $\widetilde{\eta}^{\left(n\right)}$, satisfying the “non-alternating analogue” equation $\gamma^{\left(n\right)}+\widetilde{\delta}^{\left(n\right)}/e=\widetilde{\eta}^{\left(n\right)}$, and provide results parallel to those for the original system.
2. Recurrence Relations for $\gamma^{\left(n\right)}$
2.1. Basic Recurrences
Let $X\sim F_{X}$ be an arbitrary random variable with finite $n^{\textrm{th}}$ raw moment $\textrm{E}_{X}\left[X^{n}\right]$, for $n\in\mathbb{Z}_{\geq 0}$, and $\ell^{\textrm{th}}$ cumulant $\kappa_{\ell}$, for $\ell\in\mathbb{Z}_{\geq 1}$. The recurrences
$\textrm{E}_{X}\left[X^{n}\right]={\displaystyle\sum_{j=0}^{n-1}}\binom{n-1}{j}\kappa_{n-j}\textrm{E}_{X}\left[X^{j}\right],\quad n\in\mathbb{Z}_{\geq 1}$ (9)
and
$\kappa_{\ell}=\textrm{E}_{X}\left[X^{\ell}\right]-{\displaystyle\sum_{i=0}^{\ell-2}}\binom{\ell-1}{i}\kappa_{i+1}\textrm{E}_{X}\left[X^{\ell-i-1}\right],\quad\ell\in\mathbb{Z}_{\geq 2}$ (10)
then follow from the complete exponential Bell polynomials $\textrm{E}_{X}\left[X^{n}\right]=B_{n}\left(\kappa_{1},\kappa_{2},\ldots,\kappa_{n}\right)$, with $\textrm{E}_{X}\left[X^{0}\right]=1$ and $\kappa_{1}=\textrm{E}_{X}\left[X\right]$.
For the case of $X\equiv X_{\textrm{G}}\sim\textrm{Gumbel}\left(0,1\right)$, the identities in (9) and (10) can be simplified by setting
$\kappa_{\ell}=\begin{cases}\gamma,&\ell=1\\ \left(\ell-1\right)!\zeta\left(\ell\right),&\ell\in\left\{2,3,\ldots\right\}\end{cases},$
immediately yielding the corresponding equations
$\gamma^{\left(n\right)}=\gamma\cdot\gamma^{\left(n-1\right)}+{\displaystyle\sum_{j=0}^{n-2}}\dfrac{\left(n-1\right)!}{j!}\zeta\left(n-j\right)\gamma^{\left(j\right)},\quad n\in\mathbb{Z}_{\geq 1}$ (11)
and
$\left(\ell-1\right)!\zeta\left(\ell\right)=\gamma^{\left(\ell\right)}-\gamma\cdot\gamma^{\left(\ell-1\right)}-{\displaystyle\sum_{i=1}^{\ell-2}}\dfrac{\left(\ell-1\right)!}{\left(\ell-i-1\right)!}\zeta\left(i+1\right)\gamma^{\left(\ell-i-1\right)},\quad\ell\in\mathbb{Z}_{\geq 2}.$ (12)
Further, we can rewrite (11) as:
$\begin{array}[]{c}\gamma^{\left(1\right)}=\gamma,\\ \gamma^{\left(2\right)}=\gamma^{2}+\zeta\left(2\right),\\ \gamma^{\left(3\right)}=\gamma^{3}+3\zeta\left(2\right)\gamma+2\zeta\left(3\right),\\ \gamma^{\left(4\right)}=\gamma^{4}+6\zeta\left(2\right)\gamma^{2}+8\zeta\left(3\right)\gamma+\dfrac{27}{2}\zeta\left(4\right),\\ \cdots;\end{array}$ (13)
and solve (12) recursively to obtain:
$\begin{array}[]{c}\zeta\left(2\right)=\gamma^{\left(2\right)}-\gamma^{2},\\ \zeta\left(3\right)=\dfrac{1}{2}\left(\gamma^{\left(3\right)}-3\gamma^{\left(2\right)}\gamma+2\gamma^{3}\right),\\ \zeta\left(4\right)=\dfrac{1}{6}\left[\gamma^{\left(4\right)}-4\gamma^{\left(3\right)}\gamma-3\left(\gamma^{\left(2\right)}\right)^{2}+12\gamma^{\left(2\right)}\gamma^{2}-6\gamma^{4}\right],\\ \zeta\left(5\right)=\dfrac{1}{24}\left[\gamma^{\left(5\right)}-5\gamma^{\left(4\right)}\gamma-10\gamma^{\left(3\right)}\gamma^{\left(2\right)}+20\gamma^{\left(3\right)}\gamma^{2}+30\left(\gamma^{\left(2\right)}\right)^{2}\gamma-60\gamma^{\left(2\right)}\gamma^{3}+24\gamma^{5}\right],\\ \cdots.\end{array}$ (14)
Replacing $\pi^{2}$ by $6\zeta\left(2\right)$ in the even-zeta identity
$\zeta\left(2n\right)=\dfrac{\left(-1\right)^{n+1}\mathscr{B}_{2n}}{2\left(2n\right)!}\left(2\pi\right)^{2n}$
(where $\mathscr{B}_{2n}$ denotes the $\left(2n\right)^{\textrm{th}}$ Bernoulli number) then gives
$\zeta\left(2n\right)=\dfrac{\left(-1\right)^{n+1}2^{3n-1}3^{n}\mathscr{B}_{2n}}{\left(2n\right)!}\left(\zeta\left(2\right)\right)^{n}=\dfrac{\left(-1\right)^{n+1}2^{3n-1}3^{n}\mathscr{B}_{2n}}{\left(2n\right)!}\left(\gamma^{\left(2\right)}-\gamma^{2}\right)^{n},$ (15)
and we can set (15) equal to the expression for $\zeta\left(2n\right)$ generated by the system (14) for any $n\in\mathbb{Z}_{\geq 2}$. After clearing denominators, this yields a polynomial equation of degree $2n$ in $\gamma$,
$P_{n}\left(\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\ldots,\gamma^{\left(2n\right)}\right)=\alpha_{n,0}\gamma^{\left(2n\right)}-\alpha_{n,1}\gamma^{\left(2n-1\right)}\gamma-\sum_{i=0}^{2n}\left[\sum_{j=1}^{J_{i}}\alpha_{n,i,j}{\displaystyle\prod_{k=2}^{2n-2}}\left(\gamma^{\left(k\right)}\right)^{p_{n,k,i,j}}\right]\gamma^{i}=0$ (16)
for some $J_{i}\in\mathbb{Z}_{\geq 1}$, $\alpha_{n,0},\alpha_{n,1}\in\mathbb{Z}_{\neq 0}$, $\alpha_{n,i,j}\in\mathbb{Z}$, and $p_{n,k,i,j}\in\mathbb{Z}_{\geq 0}$. In particular, for $n=2$ and $3$ we have
$P_{2}\left(\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\gamma^{\left(4\right)}\right)=5\gamma^{\left(4\right)}-20\gamma^{\left(3\right)}\gamma-27\left(\gamma^{\left(2\right)}\right)^{2}+84\gamma^{\left(2\right)}\gamma^{2}-42\gamma^{4}=0$
and
$\begin{array}[]{c}P_{3}\left(\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\gamma^{\left(4\right)},\gamma^{\left(5\right)},\gamma^{\left(6\right)}\right)=7\gamma^{\left(6\right)}-42\gamma^{\left(5\right)}\gamma+210\gamma^{\left(4\right)}\gamma^{2}-105\gamma^{\left(4\right)}\gamma^{\left(2\right)}-70\left(\gamma^{\left(3\right)}\right)^{2}\\ +840\gamma^{\left(3\right)}\gamma^{\left(2\right)}\gamma-840\gamma^{\left(3\right)}\gamma^{3}+18\left(\gamma^{\left(2\right)}\right)^{3}-1314\left(\gamma^{\left(2\right)}\right)^{2}\gamma^{2}+1944\gamma^{\left(2\right)}\gamma^{4}-648\gamma^{6}=0,\end{array}$
respectively.
2.2. Variable Elimination
For $n\in\mathbb{Z}_{\geq 2}$, define $\mathcal{G}_{n}^{\left(2n\right)}=\left\{\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\ldots,\gamma^{\left(2n\right)}\right\}$, $\mathcal{G}_{n}^{\left(\textrm{L}\right)}=\left\{\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\ldots,\gamma^{\left(n-1\right)}\right\}$, and $\mathcal{G}_{n}^{\left(\textrm{U}\right)}=\left\{\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$. Since
$\mathbb{Z}\left[\mathcal{G}_{n}^{\left(2n\right)}\setminus\left\{\gamma\right\}\right]\left[\gamma\right]=\mathbb{Z}\left[\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right]\left[\mathcal{G}_{n}^{\left(\textrm{L}\right)}\setminus\left\{\gamma\right\}\right],$
we may view each $P_{n}$ given by (16) as a polynomial in the variables $\gamma^{\left(k\right)}\in\left(\mathcal{G}_{n}^{\left(\textrm{L}\right)}\setminus\left\{\gamma\right\}\right)$ with coefficients in $\mathbb{Z}\left[\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right]$, and then transform the system
$P_{m}\left(\mathcal{G}_{m}^{\left(2m\right)}\right)=0,\quad m\in\left\{2,3,\ldots,n\right\}$
into
$\widetilde{P}_{m}\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)=0,\quad m\in\left\{2,3,\ldots,n\right\}$
by eliminating all $\gamma^{\left(k\right)}\in\left(\mathcal{G}_{m}^{\left(\textrm{L}\right)}\setminus\left\{\gamma\right\}\right)$ for each $m$. As will be shown in Section 2.2.2, this yields at least one non-zero polynomial $\widetilde{P}_{n,\theta}\in\mathbb{Z}\left[\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\right]$ for each pair $\left(n,\theta\right):n\in\mathbb{Z}_{\geq 2},\theta\in\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)$ such that
$\widetilde{P}_{n,\theta}\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)=0$ (17)
and $\widetilde{P}_{n,\theta}$ can be viewed as a non-trivial polynomial in $\theta$ (i.e., $\deg_{\theta}\left(\widetilde{P}_{n,\theta}\right)\geq 1$).
2.2.1. Algebraic Structure of the $P_{n}$
Let $\mathcal{R}_{n}=\mathbb{Q}\left[\mathcal{G}_{n}^{\left(2n\right)}\right]$ denote a polynomial ring in $2n$ variables over the rationals. For $m\in\left\{2,3,\ldots,n\right\}$, the polynomial equations $P_{m}=0$ generate an ideal $\mathcal{I}_{n}=\left\langle P_{2},P_{3},\ldots,P_{n}\right\rangle\subset\mathcal{R}_{n}$, and we let $\mathcal{V}_{n}\subset\mathbb{C}^{2n}$ denote the algebraic variety defined by this ideal. From (16), it is clear that each successive $P_{n}$ introduces the new variable $\gamma^{\left(2n\right)}$ (which does not appear in $\left\{P_{2},P_{3},\ldots,P_{n-1}\right\}$) as a linear term with non-zero integer coefficient derived from the coefficient of $\gamma^{\left(2n\right)}$ in the $\zeta\left(2n\right)$ expansion of (14).
Lemma 1.
The set of polynomials $\left\{P_{2},P_{3},\ldots,P_{n}\right\}$ are algebraically independent.
?proofname? .
Assume (for purposes of contradiction) that the polynomials $\left\{P_{2},P_{3},\ldots,P_{n}\right\}$ are algebraically dependent; that is, there exists a non-zero polynomial $Q_{n}\in\mathbb{Q}\left[t_{2},t_{3},\ldots,t_{n}\right]$ such that $Q_{n}\left(P_{2},P_{3},\ldots,P_{n}\right)=0$. Writing $Q_{n}$ as a polynomial in $t_{n}$ gives
$Q_{n}=\sum_{d=0}^{D_{n}}C_{d}^{\left(n\right)}\left(t_{2},t_{3},\ldots,t_{n-1}\right)t_{n}^{d}$
for some $D_{n}\in\mathbb{Z}_{\geq 1}$ and $C_{d}^{\left(n\right)}\in\mathbb{Q}\left[t_{2},t_{3},\ldots,t_{n-1}\right]$, with $C_{D_{n}}^{\left(n\right)}\not\equiv 0$, and we can substitute $P_{m}$ for $t_{m}$ for all $m\in\left\{2,3,\ldots,n\right\}$, obtaining
$\sum_{d=0}^{D_{n}}C_{d}^{\left(n\right)}\left(P_{2},P_{3},\ldots,P_{n-1}\right)P_{n}^{d}=0.$ (18)
Then, since
$P_{n}=\alpha_{n,0}\gamma^{\left(2n\right)}+U_{n}\left(\mathcal{G}_{n}^{\left(2n\right)}\setminus\left\{\gamma^{\left(2n\right)}\right\}\right),$
where $\alpha_{n,0}\in\mathbb{Z}_{\neq 0}$, $U_{n}$ is defined implicitly by (16), and the polynomials $\left\{P_{2},P_{3},\ldots,P_{n-1}\right\}$ are independent of $\gamma^{\left(2n\right)}$, it follows that the highest-degree term in $\gamma^{\left(2n\right)}$ within the expansion of $Q_{n}\left(P_{2},P_{3},\ldots,P_{n}\right)$ is $C_{D_{n}}^{\left(n\right)}\left(P_{2},P_{3},\ldots,P_{n-1}\right)\left(\alpha_{n,0}\gamma^{\left(2n\right)}\right)^{D_{n}}$.
To satisfy (18), this highest-degree term must equal 0 (because the $\left\{P_{2},P_{3},\ldots,P_{n-1}\right\}$ are independent of $\gamma^{\left(2n\right)}$), forcing $C_{D_{n}}^{\left(n\right)}\left(P_{2},P_{3},\ldots,P_{n-1}\right)=0$ and the polynomials $\left\{P_{2},P_{3},\ldots,P_{n-1}\right\}$ to be algebraically dependent. Similarly, it can be shown by iteration that $C_{D_{m+1}}^{\left(m+1\right)}\left(P_{2},P_{3},\ldots,P_{m}\right)=0$ for all $m\in\left\{2,3,\ldots,n-1\right\}$. For $m=2$, this means there exists a non-zero polynomial $C_{D_{3}}^{\left(3\right)}\left(t_{2}\right)\in\mathbb{Q}\left[t_{2}\right]$ in $\mathcal{R}_{n}$ such that $C_{D_{3}}^{\left(3\right)}\left(P_{2}\right)=0$. Writing $C_{D_{3}}^{\left(3\right)}\left(t_{2}\right)$ as
$C_{D_{3}}^{\left(3\right)}\left(t_{2}\right)={\displaystyle\sum_{d^{\prime}=0}^{D^{\prime}}}c_{d^{\prime}}t_{2}^{d^{\prime}}$
for some $D^{\prime}\in\mathbb{Z}_{\geq 1}$ and $c_{d^{\prime}}\in\mathbb{Q}$, with $c_{D^{\prime}}\neq 0$, we then see that
$C_{D_{3}}^{\left(3\right)}\left(P_{2}\right)={\displaystyle\sum_{d^{\prime}=0}^{D^{\prime}}}c_{d^{\prime}}\left(5\gamma^{\left(4\right)}+U_{2}\left(\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)}\right)\right)^{d^{\prime}}$
is a polynomial in the variable $\gamma^{\left(4\right)}$ with leading coefficient $c_{D^{\prime}}\left(5\right)^{D^{\prime}}\neq 0$. Thus, $C_{D_{3}}^{\left(3\right)}\left(P_{2}\right)$ is a non-zero element of the integral domain $\mathcal{R}_{n}$ – a clear contradiction – implying that the set $\left\{P_{2},P_{3},\ldots,P_{n}\right\}$ must be algebraically independent. ∎
Lemma 2.
The variety $\mathcal{V}_{n}$ is irreducible and has function field of transcendence degree $n+1$.
?proofname?.
To determine the dimension of the variety $\mathcal{V}_{n}$, we exploit the triangular structure of the polynomials given by (16). For each $m\in\left\{2,3,\ldots,n\right\}$, one can write
$P_{m}=\alpha_{m,0}\gamma^{\left(2m\right)}+U_{m}\left(\mathcal{G}_{m}^{\left(2m\right)}\setminus\left\{\gamma^{\left(2m\right)}\right\}\right),\quad\alpha_{m,0}\in\mathbb{Z}_{\neq 0},$
where $\gamma^{\left(2m\right)}$ does not occur in $\mathcal{I}_{m-1}$. Since $\alpha_{m,0}\in\mathbb{Q}^{\times}$, the relation $P_{m}=0$ identifies $\gamma^{\left(2m\right)}+\alpha_{m,0}^{-1}U_{m}=0$ in the quotient ring $\mathcal{R}_{n}/\mathcal{I}_{m}$, yielding
$\mathcal{R}_{n}/\mathcal{I}_{m}\cong\mathbb{Q}\left[\mathcal{G}_{n}^{\left(2n\right)}\setminus\left\{\gamma^{\left(4\right)},\gamma^{\left(6\right)},\ldots,\gamma^{\left(2m\right)}\right\}\right].$
Iterating from $m=2$ then gives
$\mathcal{R}_{n}/\mathcal{I}_{m-1}\cong\mathbb{Q}\left[\mathcal{G}_{n}^{\left(2n\right)}\setminus\left\{\gamma^{\left(4\right)},\gamma^{\left(6\right)},\ldots,\gamma^{\left(2m-2\right)}\right\}\right],$
which is a polynomial ring over $\mathbb{Q}$, and thus an integral domain.
It then follows that $P_{m}\not\equiv 0$ in $\mathcal{R}_{n}/\mathcal{I}_{m-1}$ (because $P_{m}$ contains $\gamma^{\left(2m\right)}$ with a non-zero coefficient, implying that $\gamma^{\left(2m\right)}$ remains one of the polynomial variables in the ring $\mathcal{R}_{n}/\mathcal{I}_{m-1}$), and the polynomials $P_{2},P_{3},\ldots,P_{n}$ form a regular sequence of length n1n-1 in $\mathcal{R}_{n}$. Since $\mathcal{R}_{n}$ is a polynomial ring over $\mathbb{Q}$, it must be Cohen-Macaulay, making $\mathcal{I}_{n}$ a complete intersection of height n1n-1 and
$\dim\left(\mathcal{V}_{n}\right)=\dim\left(\mathcal{R}_{n}\right)-\textrm{ht}\left(\mathcal{I}_{n}\right)=2n-\left(n-1\right)=n+1.$
Finally, since
$\mathcal{R}_{n}/\mathcal{I}_{n}\cong\mathbb{Q}\left[\mathcal{G}_{n}^{\left(2n\right)}\setminus\left\{\gamma^{\left(4\right)},\gamma^{\left(6\right)},\ldots,\gamma^{\left(2n\right)}\right\}\right]$
is a polynomial ring (and thus an integral domain), $\mathcal{I}_{n}$ must be prime and $\mathcal{V}_{n}$ irreducible, confirming that $\mathcal{V}_{n}$ has a function field of transcendence degree $n+1$. ∎
2.2.2. Existence and Non-Triviality of $\widetilde{P}_{n,\theta}=0$
The following two lemmas demonstrate the existence of a polynomial $\widetilde{P}_{n,\theta}\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)$ asserted in (17).
Lemma 3.
For any $n\in\mathbb{Z}_{\geq 2}$, there exists at least one non-zero irreducible polynomial $\widetilde{P}_{n}\in\mathbb{Z}\left[\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right]$ such that $\widetilde{P}_{n}\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)=0$.
?proofname?.
We know from Lemma 2 that the variables in $\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}$, as elements of $\mathcal{V}_{n}$, belong to a function field of transcendence degree $n+1$. Then, since $\left|\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right|=n+2$, it follows that these variables must be algebraically dependent on $\mathcal{V}_{n}$, implying that there exists at least one non-zero irreducible polynomial relation among them; that is, $\widetilde{P}_{n}\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)=0$ with $\widetilde{P}_{n}\in\mathbb{Z}\left[\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right]$. ∎
Lemma 4.
For any pair $\left(n,\theta\right):n\in\mathbb{Z}_{\geq 2},\theta\in\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)$, there exists at least one non-zero polynomial $\widetilde{P}_{n,\theta}\in\mathbb{Z}\left[\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right]\left[\theta\right]$ with $\deg_{\theta}\left(\widetilde{P}_{n,\theta}\right)\geq 1$ such that $\widetilde{P}_{n,\theta}\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)=0$.
?proofname?.
To confirm the existence of a polynomial equation expressed in terms of a particular element $\theta\in\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)$, it suffices to show that the remaining $n+1$ variables, $\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\},$ are algebraically independent on $\mathcal{V}_{n}$. In that case, $\theta$ is algebraic over the other $n+1$ variables, and thus satisfies a non-trivial polynomial equation
$\widetilde{P}_{n,\theta}\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)=\sum_{d=0}^{D_{n,\theta}}A_{d}^{\left(n,\theta\right)}\left(\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right)\theta^{d}=0$ (19)
for some $D_{n,\theta}\in\mathbb{Z}_{\geq 1}$ and $A_{d}^{\left(n,\theta\right)}\in\mathbb{Z}\left[\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right]$, with $A_{D_{n,\theta}}^{\left(n,\theta\right)}\not\equiv 0$.
To assess the posited algebraic independence, we consider the algebraic structure of the polynomial $P_{n}$ in (16) in terms of relationships between moments $\gamma^{\left(\bullet\right)}$ and cumulants $\kappa_{\bullet}$ of the $\textrm{Gumbel}(0,1)$ distribution. Essentially, the constraints defining $\mathcal{V}_{n}$ (i.e., the equations $P_{m}=0,\>m\in\left\{2,3,\ldots,n\right\}$) are polynomial realizations of the identities derived from the even-zeta formula of (15). These identities imply that the even cumulants, $\kappa_{2m}=\left(2m-1\right)!\zeta\left(2m\right)$, are rationally proportional to powers of $\kappa_{2}=\zeta\left(2\right)$; that is,
$\kappa_{2m}=b_{m}\kappa_{2}^{m},\quad m=2,3,\ldots,n,$ (20)
where $b_{m}\in\mathbb{Q}^{\times}$. Furthermore, these n1n-1 constraints allow us to parameterize the variety $\mathcal{V}_{n}$ using the set of $n+1$ unconstrained algebraically independent cumulants,
$\mathcal{K}_{n}=\left\{\kappa_{1},\kappa_{2},\kappa_{3},\kappa_{5},\ldots,\kappa_{2n-1}\right\}.$
In this way, the $\gamma^{\left(j\right)}$, for $j\in\left\{1,2,\ldots,2n\right\}$, are polynomials in the full set of $\kappa_{\ell}$, and can be viewed as functions of the components of $\mathcal{K}_{n}$ via (9).
The moment-cumulant correspondence (truncated at $2n$) defines a polynomial change of variables on $\mathbb{A}^{2n}$ (the affine space of all $2n$-tuples), and this transformation is invertible with polynomial inverse given by the classical truncated moment-cumulant inversion formulas (up to order $2n$). Under this change of variables, the defining relations $P_{m}=0,\>m\in\left\{2,3,\ldots,n\right\}$ for $\mathcal{V}_{n}$ are precisely the images of the cumulant constraints of (20), and all polynomial relations (e.g., algebraic independence) among the $\gamma^{\left(j\right)}$ on $\mathcal{V}_{n}$ can be verified in cumulant coordinates. Consequently, we may test for algebraic independence via the Jacobian computations below, carried out in the components of $\mathcal{K}_{n}$.
According to the Jacobian criterion for algebraic independence (which is applicable because we are working over a field of characteristic 0), the $n+1$ polynomials corresponding to $\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}$, for any $\theta\in\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)$, are algebraically independent if and only if the determinant of the $\left(n+1\right)\times\left(n+1\right)$ Jacobian matrix
$\mathbf{J}_{-\theta}^{\left(n\right)}\left(j,\ell\right)=\left[\frac{d\gamma^{\left(j\right)}}{d\kappa_{\ell}}\right],\quad j\in\left(\left\{1,n,n+1,\ldots,2n\right\}\setminus\left\{k:\theta=\gamma^{\left(k\right)}\right\}\right),\ell\in\left\{1,2,3,5,\ldots,2n-1\right\}$
is not identically zero. To determine whether or not this condition holds, first note that
$\frac{d\gamma^{\left(j\right)}}{d\kappa_{\ell}}=\frac{\partial\gamma^{\left(j\right)}}{\partial\kappa_{\ell}}+\sum_{m=2}^{n}\frac{\partial\gamma^{\left(j\right)}}{\partial\kappa_{2m}}\frac{d\kappa_{2m}}{d\kappa_{\ell}},$
which simplifies to
$\frac{d\gamma^{\left(j\right)}}{d\kappa_{\ell}}=\begin{cases}{\displaystyle\frac{\partial\gamma^{\left(j\right)}}{\partial\kappa_{2}}+\sum_{m=2}^{n}\frac{\partial\gamma^{\left(j\right)}}{\partial\kappa_{2m}}mb_{m}\kappa_{2}^{m-1}},&\ell=2\\ {\displaystyle\frac{\partial\gamma^{\left(j\right)}}{\partial\kappa_{\ell}}},&\ell\in\left\{1,3,5,\ldots,2n-1\right\}\end{cases}$
because $d\kappa_{2m}/d\kappa_{2}=mb_{m}\kappa_{2}^{m-1}$ from (20) and $d\kappa_{2m}/d\kappa_{\ell}\equiv 0$ for all $\ell\neq 2$. Specializing the Jacobian at the convenient realization $\mathcal{K}_{n}^{*}=\left\{\kappa_{1}=1,\kappa_{2}=0,\kappa_{3}=0,\kappa_{5}=0,\ldots,\kappa_{2n-1}=0\right\}$, for which $mb_{m}\kappa_{2}^{m-1}=0$, then gives
$\left.\mathbf{J}_{-\theta}^{\left(n\right)}\left(j,\ell\right)\right|_{\mathcal{K}_{n}^{*}}=\left.\left[\frac{\partial\gamma^{\left(j\right)}}{\partial\kappa_{\ell}}\right]\right|_{\mathcal{K}_{n}^{*}},$
which simplifies further to
$\left.\mathbf{J}_{-\theta}^{\left(n\right)}\left(j,\ell\right)\right|_{\mathcal{K}_{n}^{*}}=\binom{j}{\ell}$
because $\partial\gamma^{\left(j\right)}/\partial\kappa_{\ell}=\binom{j}{\ell}\gamma^{\left(j-\ell\right)}$ (from the Bell-polynomial identity $\partial B_{j}\left(\kappa_{1},\kappa_{2},\ldots,\kappa_{j}\right)=$ $\binom{j}{\ell}B_{j-\ell}\left(\kappa_{1},\kappa_{2},\ldots,\kappa_{j-\ell}\right)$) and $\gamma^{\left(j-\ell\right)}=\kappa_{1}^{j-\ell}=1$ at $\mathcal{K}_{n}^{*}$.
Given this form of the Jacobian, we can write the determinant as
$\det\left(\left.\mathbf{J}_{-\theta}^{\left(n\right)}\right|_{\mathcal{K}_{n}^{*}}\right)=\det\left(\mathbf{M}_{-\theta}^{\left(n\right)}\right),$
where $\mathbf{M}_{-\theta}^{\left(n\right)}$ is the $\left(n+1\right)\times\left(n+1\right)$ matrix
$\mathbf{M}_{-\theta}^{\left(n\right)}\left(j,\ell\right)=\binom{j}{\ell},\quad j\in\left(\left\{1,n,n+1,\ldots,2n\right\}\setminus\left\{k:\theta=\gamma^{\left(k\right)}\right\}\right),\ell\in\left\{1,2,3,5,\ldots,2n-1\right\}.$
To see that $\det\left(\mathbf{M}_{-\theta}^{\left(n\right)}\right)\neq 0$, we note that $\mathbf{M}_{-\theta}^{\left(n\right)}$ is a submatrix of the lower-triangular Pascal matrix,
$\mathbf{P}^{\left(n\right)}\left(r,c\right)=\binom{r}{c},\quad r\in\left\{0,1,\ldots,2n\right\},c\in\left\{0,1,\ldots,2n\right\},$
which satisfies the conditions for a lower strictly totally positive (LSTP) matrix given by Section 2.3, Theorem 2.8 of Pinkus [8]. Letting $j_{i}$ and $\ell_{i}$ denote, respectively, the $i^{\textrm{th}}$ row and column indices of the submatrix $\mathbf{M}_{-\theta}^{\left(n\right)}$, it then follows from the definition of an LSTP matrix (see paragraph 2, Section 2.3 of Pinkus [8]) that if $j_{i}\geq\ell_{i}$ for all $i\in\left\{1,2,\ldots,n+1\right\}$, then $\det\left(\mathbf{M}_{-\theta}^{\left(n\right)}\right)>0$.
The condition $j_{i}\geq\ell_{i}$ can be verified for all possible choices of $\theta\in\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)$ (for $n\in\mathbb{Z}_{\geq 2}$) through cases (I)-(V) below, confirming that $\det\left(\mathbf{M}_{-\theta}^{\left(n\right)}\right)\neq 0\Longrightarrow\det\left(\left.\mathbf{J}_{-\theta}^{\left(n\right)}\right|_{\mathcal{K}_{n}^{*}}\right)\neq 0\Longrightarrow\det\left(\mathbf{J}_{-\theta}^{\left(n\right)}\right)\not\equiv 0$, and thus that the set of elements $\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}$ are algebraically independent on $\mathcal{V}_{n}$ for all such $\theta$.
(I) $\theta=\gamma^{\left(2n\right)}$. The row indices are $\left\{1,n,n+1,\ldots,2n-1\right\}$, and by inspection, $j_{1}=\ell_{1}$ and $j_{2}\geq\ell_{2}$. Then, for $i\in\left\{3,4,\ldots,n+1\right\}$, we have $j_{i}=n-2+i$ and $\ell_{i}=2i-3$, so $j_{i}\geq\ell_{i}$ for all relevant $i$.
(II) $\theta=\gamma$. The row indices are $\left\{n,n+1,\ldots,2n\right\}$, and by inspection, $j_{1}>\ell_{1}$ and $j_{2}>\ell_{2}$. For $i\in\left\{3,4,\ldots,n+1\right\}$, we have $j_{i}=n-1+i$ and $\ell_{i}=2i-3$, so $j_{i}\geq\ell_{i}$ for all relevant $i$.
(III) $\theta=\gamma^{\left(n\right)}$. The row indices are $\left\{1,n+1,n+2,\ldots,2n\right\}$, and by inspection, $j_{1}=\ell_{1}$ and $j_{2}>\ell_{2}$. Then, for $i\in\left\{3,4,\ldots,n+1\right\}$, we have $j_{i}=n-1+i$ and $\ell_{i}=2i-3$, so $j_{i}\geq\ell_{i}$ as in case (II).
(IV) $\theta=\gamma^{\left(n+1\right)}$. The row indices are $\left\{1,n,n+2,n+3,\ldots,2n\right\}$, and by inspection, $j_{1}=\ell_{1}$ and $j_{2}\geq\ell_{2}$. Then, for $i\in\left\{3,4,\ldots,n+1\right\}$, we have $j_{i}=n-1+i$ and $\ell_{i}=2i-3$, so $j_{i}\geq\ell_{i}$ as in case (II).
(V) $\theta=\gamma^{\left(k\right)}$, where $k\in\left\{n+2,n+3,\ldots,2n-1\right\}$. The row indices are $\left\{1,n,n+1,\ldots,k-1,k+1,\ldots,2n\right\}$, with k1k-1 appearing as the kn+1k-n+1 index, and $k+1$ as the kn+2k-n+2 index. By inspection, $j_{1}=\ell_{1}$ and $j_{2}\geq\ell_{2}$. Then, for $i\in\left\{3,4,\ldots,k-n+1\right\}$, we have $j_{i}=n-2+i$ and $\ell_{i}=2i-3$, so $j_{i}\geq\ell_{i}$ as in case (I); and for $i\in\left\{k-n+2,k-n+3,\ldots,n+1\right\}$, we have $j_{i}=n-1+i$ and $\ell_{i}=2i-3$, so $j_{i}\geq\ell_{i}$ as in case (II). ∎
2.3. Transcendence Results
The transformation of the original polynomials of (16) into the reduced form of (17) lays the foundation for part (ii) of Theorem 1 below. However, one additional condition on the $\widetilde{P}_{n,\theta}$ is required. Although we know there exists $\widetilde{P}_{n,\theta}\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)=0$ that can be expressed as a non-trivial polynomial equation in $\theta\in\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)$ of the form given by (19), it remains possible that $A_{d}^{\left(n,\theta\right)}=0$ for all $d\in\left\{1,2,\ldots,D_{n,\theta}\right\}$ at the actual realizations of the elements $\gamma^{\left(k\right)}\in\mathcal{G}_{n}^{\left(2n\right)}$ (to be denoted by $\boldsymbol{\gamma^{*}}\Longleftrightarrow\gamma^{\left(k\right)}=\textrm{E}_{X}\left[X_{\textrm{G}}^{k}\right]\>\forall\>k\in\mathbb{Z}_{\geq 1}$). Lemma 5 below, supported by Conjecture 1, ensures this does not happen.
Recalling $\mathcal{R}_{n}=\mathbb{Q}\left[\mathcal{G}_{n}^{\left(2n\right)}\right]$ and $\mathcal{I}_{n}=\left\langle P_{2},P_{3},\ldots,P_{n}\right\rangle\subset\mathcal{R}_{n}$ from Section 2.2.1, let $\mathcal{S}_{n}=\mathcal{R}_{n}/\mathcal{I}_{n}$ and define $\mathcal{J}_{n,\boldsymbol{\gamma^{*}}}\subset\mathcal{R}_{n}$ as the kernel of the evaluation homomorphism induced by $\boldsymbol{\gamma^{*}}$ on $\mathcal{R}_{n}$. Since
$\left.P_{m}\left(\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\ldots,\gamma^{\left(2m\right)}\right)\right|_{\boldsymbol{\gamma^{*}}}=0$
for all $m\in\left\{2,3,\ldots,n\right\}$, we then have $\mathcal{I}_{n}\subset\mathcal{J}_{n,\boldsymbol{\gamma^{*}}}$, implying that $\mathcal{\overline{J}}_{n,\boldsymbol{\gamma^{*}}}=\mathcal{J}_{n,\boldsymbol{\gamma^{*}}}/\mathcal{I}_{n}$ is a prime ideal of $\mathcal{S}_{n}$. Now define the $\left(m-1\right)\times\left(m-1\right)$ Jacobian matrix
$\mathbf{J}_{m,\theta}^{\left(n\right)}\left(i,j\right)=\left[\frac{\partial P_{i}}{\partial\gamma^{\left(j\right)}}\right],\quad i\in\left\{2,3,\ldots,m\right\},j\in\left(\left\{2,3,\ldots,m-1\right\}\cup\left\{k:\theta=\gamma^{\left(k\right)}\right\}\right),$
for which $\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)\in\mathcal{R}_{n}.$
Lemma 5.
For fixed $n\in\mathbb{Z}_{\geq 2}$, $m\in\left\{2,3,\ldots,n\right\}$, and $\theta\in\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)$, if $\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)\notin\mathcal{J}_{n,\boldsymbol{\gamma^{*}}}$ – or equivalently, if
$\left.\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)\right|_{\boldsymbol{\gamma^{*}}}\neq 0$ (21)
– then there exists a polynomial $\widetilde{P}_{m,\theta}\in\mathbb{Z}\left[\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right]$ satisfying (17) such that, when written in the form of (19), at least one coefficient $A_{d}^{\left(m,\theta\right)}\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right),\>d\geq 1$ must remain non-zero at $\boldsymbol{\gamma^{*}}$.
?proofname?.
For given $n\in\mathbb{Z}_{\geq 2}$, let $\mathcal{R}_{m}=\mathbb{Q}\left[\mathcal{G}_{m}^{\left(2m\right)}\right]\subset\mathcal{R}_{n}$, $\mathcal{I}_{m}=\left\langle P_{2},P_{3},\ldots,P_{m}\right\rangle\subset\mathcal{R}_{m}$, and $\mathcal{S}_{m}=\mathcal{R}_{m}/\mathcal{I}_{m}$ for $m\in\left\{2,3,\ldots,n\right\}$, and define $\mathcal{J}_{m,\boldsymbol{\gamma^{*}}}=\mathcal{J}_{n,\boldsymbol{\gamma^{*}}}\cap\mathcal{R}_{m}$ and $\mathcal{\overline{J}}_{m,\boldsymbol{\gamma^{*}}}=\mathcal{J}_{m,\boldsymbol{\gamma^{*}}}/\mathcal{I}_{m}$, so that $\mathcal{\overline{J}}_{m,\boldsymbol{\gamma^{*}}}$ is a prime ideal of $\mathcal{S}_{m}$. Since $\mathcal{I}_{m}\subset\mathcal{J}_{m,\boldsymbol{\gamma^{*}}}$, the specialization at $\boldsymbol{\gamma^{*}}$ induces a well-defined homomorphism $\mathcal{S}_{m}=\mathcal{R}_{m}/\mathcal{I}_{m}\rightarrow\mathbb{C}$ whose kernel is $\mathcal{\overline{J}}_{m,\boldsymbol{\gamma^{*}}}$; in particular, $s\notin\mathcal{\overline{J}}_{m,\boldsymbol{\gamma^{*}}}$ implies $\left.s\right|_{\boldsymbol{\gamma^{*}}}\neq 0$.
Now let $\mathcal{U}_{m,\theta}=\mathbb{Q}\left[\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right]$ for $\theta\in\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)$, and define
$\varphi:\mathcal{U}_{m,\theta}\rightarrow\mathcal{S}_{m}$
as the natural homomorphism sending each generator of $\mathcal{U}_{m,\theta}$ to its class in $\mathcal{S}_{m}$, with contraction $\mathcal{T}_{m,\theta}=\varphi^{-1}\left(\mathcal{\overline{J}}_{m,\boldsymbol{\gamma^{*}}}\right)\subset\mathcal{U}_{m,\theta}$. By construction, an element $u\in\mathcal{U}_{m,\theta}$ satisfies $u\notin\mathcal{T}_{m,\theta}$ if and only if $u$ does not vanish at $\boldsymbol{\gamma^{*}}$. Thus, whenever $u\notin\mathcal{T}_{m,\theta}$, the specialization $\boldsymbol{\gamma^{*}}$ extends to the localization $\left(\mathcal{U}_{m,\theta}\right)_{u}$.
Recognizing that $\mathcal{R}_{m}$ is (non-canonically) isomorphic to a polynomial ring $\mathcal{U}_{m,\theta}\left[\gamma^{\left(2\right)},\gamma^{\left(3\right)},\ldots,\gamma^{\left(m-1\right)},\theta\right]$, we see that $\mathcal{S}_{m}$ is a finitely presented $\mathcal{U}_{m,\theta}$-algebra with generators $\left\{\gamma^{\left(2\right)},\gamma^{\left(3\right)},\ldots,\gamma^{\left(m-1\right)},\theta\right\}$ and defining relations $\left\{P_{2},P_{3},\ldots,P_{m}\right\}$. In particular, $\varphi:\mathcal{U}_{m,\theta}\rightarrow\mathcal{S}_{m}$ corresponds to a morphism of affine schemes
$\psi:\textrm{Spec}\left(\mathcal{S}_{m}\right)\rightarrow\textrm{Spec}\left(\mathcal{U}_{m,\theta}\right)$
of finite presentation (and thus of finite type), where $\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)$ is the Jacobian determinant of this presentation with respect to the variables $\left\{\gamma^{\left(2\right)},\gamma^{\left(3\right)},\ldots,\gamma^{\left(m-1\right)},\theta\right\}$. It then follows that the hypothesis $\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)\notin\mathcal{J}_{n,\boldsymbol{\gamma^{*}}}$ implies that the image of $\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)$ is not contained in $\mathcal{\overline{J}}_{m,\boldsymbol{\gamma^{*}}}$, and equivalently, $\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)$ is invertible in the local ring $\left(\mathcal{S}_{m}\right)_{\mathcal{\overline{J}}_{m,\boldsymbol{\gamma^{*}}}}$.
By the Jacobian criterion for étaleness of finitely presented morphisms in characteristic 0 (see, e.g., Sections 4.2.2 and 4.3.2 of Liu [9]), we know that $\varphi$ (and equivalently, $\psi$) is étale at $\mathcal{\overline{J}}_{m,\boldsymbol{\gamma^{*}}}$. Moreover, given that the étale locus is open and étale morphisms are locally quasi-finite, it follows (possibly after shrinking to principal open neighborhoods in both source and target) that $\psi$ is quasi-finite in a neighborhood of that point. Since $\textrm{Spec}\left(\mathcal{S}_{m}\right)\rightarrow\textrm{Spec}\left(\mathcal{U}_{m,\theta}\right)$ is affine (and thus separated) and of finite type, Zariski’s main theorem for quasi-finite separated morphisms (see, e.g., Section 4.4 of Liu [9]) then implies that there exist $u\in\mathcal{U}_{m,\theta}\setminus\mathcal{T}_{m,\theta}$ and $s\in\mathcal{S}_{m}\setminus\mathcal{\overline{J}}_{m,\boldsymbol{\gamma^{*}}}$ such that the localized $\left(\mathcal{U}_{m,\theta}\right)_{u}$-algebra $\left(\mathcal{S}_{m}\right)_{s}$ is a finite $\left(\mathcal{U}_{m,\theta}\right)_{u}$-module. (Note that $s\notin\mathcal{\overline{J}}_{m,\boldsymbol{\gamma^{*}}}$ guarantees that $s$ does not vanish at $\boldsymbol{\gamma^{*}}$, so the specialization also extends to $\left(\mathcal{S}_{m}\right)_{s}$.)
Given that $\left(\mathcal{S}_{m}\right)_{s}$ is finite over $\left(\mathcal{U}_{m,\theta}\right)_{u}$, it is clear that every element of $(\mathcal{S}_{m})_{s}$ – and in particular, the class of $\theta$ – is integral over $\left(\mathcal{U}_{m,\theta}\right)_{u}$. Thus, there exists a monic polynomial
$\begin{array}[]{c}M_{m,\theta}\left(t\right)=t^{D}+b_{D-1}^{\left(m,\theta\right)}\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right)t^{D-1}+\cdots\\ +b_{0}^{\left(m,\theta\right)}\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right)\in\left(\mathcal{U}_{m,\theta}\right)_{u}\left[t\right],\end{array}$
with $D\geq 1$ such that $M_{m,\theta}\left(\theta\right)=0$ in $\left(\mathcal{S}_{m}\right)_{s}$. Rewriting each $b_{d}^{\left(m,\theta\right)}\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right)$ as $\beta_{d}^{\left(m,\theta\right)}\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right)/u^{q_{d}^{\left(m,\theta\right)}}$ with $\beta_{d}^{\left(m,\theta\right)}\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right)\in\mathcal{U}_{m,\theta}$ and $q_{d}^{\left(m,\theta\right)}\in\mathbb{Z}_{\geq 0}$, we then can multiply $M_{m,\theta}\left(t\right)$ by $u^{\widehat{q}^{\left(m,\theta\right)}}$, where $\widehat{q}^{\left(m,\theta\right)}=\underset{d}{\max}\left\{q_{d}^{\left(m,\theta\right)}\right\}$, to obtain
$\begin{array}[]{c}M_{m,\theta}^{*}\left(t\right)=u^{\widehat{q}^{\left(m,\theta\right)}}t^{D}+\beta_{D-1}^{\left(m,\theta\right)}u^{\left(\widehat{q}^{\left(m,\theta\right)}-q_{D-1}^{\left(m,\theta\right)}\right)}\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right)t^{D-1}+\cdots\\ +\beta_{0}^{\left(m,\theta\right)}u^{\left(\widehat{q}^{\left(m,\theta\right)}-q_{0}^{\left(m,\theta\right)}\right)}\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right)\in\mathcal{U}_{m,\theta}\left[t\right],\end{array}$
with $M_{m,\theta}^{*}\left(\theta\right)=0$ in $\left(\mathcal{S}_{m}\right)_{s}$ and leading coefficient $u^{\widehat{q}^{\left(m,\theta\right)}}$. Since this equality holds in the localization $\left(\mathcal{S}_{m}\right)_{s}$, there exists $q\in\mathbb{Z}_{\geq 0}$ such that $s^{q}M_{m,\theta}^{*}\left(\theta\right)=0$ in $\mathcal{S}_{m}$.
Now specialize at $\boldsymbol{\gamma^{*}}$, and note that $u\notin\mathcal{T}_{m,\theta}$ and $s\notin\mathcal{\overline{J}}_{m,\boldsymbol{\gamma^{*}}}$ together force $\left.u\right|_{\boldsymbol{\gamma^{*}}}\neq 0$ and $\left.s\right|_{\boldsymbol{\gamma^{*}}}\neq 0$. This allows us to conclude two things. First, we see that
$\left.\left(s^{q}M_{m,\theta}^{*}\left(\theta\right)\right)\right|_{\boldsymbol{\gamma^{*}}}=0\Longrightarrow\left(\left.s\right|_{\boldsymbol{\gamma^{*}}}\right)^{q}M_{m,\theta}^{*}\left(\left.\theta\right|_{\boldsymbol{\gamma^{*}}}\right)=0$
$\Longrightarrow M_{m,\theta}^{*}\left(\left.\theta\right|_{\boldsymbol{\gamma^{*}}}\right)=0;$
and second, for the leading coefficient of $M_{m,\theta}^{*}\left(t\right)$, we find
$\left.u^{\widehat{q}^{\left(m,\theta\right)}}\right|_{\boldsymbol{\gamma^{*}}}=\left(\left.u\right|_{\boldsymbol{\gamma^{*}}}\right)^{\widehat{q}^{\left(m,\theta\right)}}\neq 0.$
These facts imply that $M_{m,\theta}^{*}\left(t\right)$ specializes to a non-zero polynomial in $t$ that vanishes at $t=\left.\theta\right|_{\boldsymbol{\gamma^{*}}}$, and so cannot be a constant. Thus, at least one coefficient of the polynomial, for degree $d\geq 1$, must remain non-zero at $\boldsymbol{\gamma^{*}}$.
Finally, let $\widehat{M}_{m,\theta}^{*}\left(t\right)$ be a primitive integer multiple of $M_{m,\theta}^{*}\left(t\right)$ (viewed in $\mathbb{Z}\left[\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right]\left[t\right]$), and factor $\widehat{M}_{m,\theta}^{*}\left(t\right)$ into irreducibles in this unique factorization domain. We then can choose an irreducible factor $\widetilde{P}_{m,\theta}\left(t\right)$ such that $\widetilde{P}_{m,\theta}\left(\left.\theta\right|_{\boldsymbol{\gamma^{*}}}\right)=0$ and $\left.\widetilde{P}_{m,\theta}\left(t\right)\right|_{\boldsymbol{\gamma^{*}}}\not\equiv 0$ (because $\left.\widehat{M}_{m,\theta}^{*}\left(t\right)\right|_{\boldsymbol{\gamma^{*}}}$ is a non-zero polynomial with root $\left.\theta\right|_{\boldsymbol{\gamma^{*}}}$). This $\widetilde{P}_{m,\theta}\in\mathbb{Z}\left[\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right]$ not only provides the non-trivial polynomial relation in $\theta$ required by (17), but also has the property that when written as a polynomial in $\theta$, it possesses at least one coefficient, for degree $d\geq 1$, that remains non-zero at $\boldsymbol{\gamma^{*}}$. ∎
The conjecture below, which we have verified computationally for all $n\in\left\{2,3,\ldots,26\right\}$, asserts that condition (21) of Lemma 5 is true. [^8] For each $n$, we executed a Python program (on an Apple M4 Max MacBook computer with 48 GB memory) to provide rigorous certification of the sufficient condition in (21). The program first evaluated $\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)$ symbolically for each $m\in\left\{2,3,\ldots,n\right\}$, and then confirmed $\left.\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)\right|_{\boldsymbol{\gamma^{*}}}\neq 0$ using interval arithmetic. The maximum value of $n$ tested (i.e., $26$) was simply the largest $n$ for which the program completed all checks within approximately 24 hours. Clearly, greater values of $n$ could be tested, but the program’s runtime increases approximately exponentially in $n$. (For example, $n=30$ would require a runtime of about 1 week.) In addition to computational verification, the conjecture is realistic on theoretical grounds because it holds for any given $n$ unless $\boldsymbol{\gamma^{*}}$ is forced to lie on one of finitely many special hypersurfaces (i.e., the vanishing loci of the minors $\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)$).
Conjecture 1.
For fixed $n\in\mathbb{Z}_{\geq 2}$,
$\left.\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)\right|_{\boldsymbol{\gamma^{*}}}\neq 0$
for all $m\in\left\{2,3,\ldots,n\right\}$ and $\theta\in\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)$.
The following theorem now provides the main results of this section.
Theorem 1.
For the sequence $\left\{\gamma^{\left(n\right)}\right\}_{n\geq 1}$:
(i) at least one element of the pair $\left\{\gamma,\gamma^{\left(2\right)}\right\}$ is transcendental;
(ii) if Conjecture 1 holds for a given $n\in\mathbb{Z}_{\geq 2}$, then at least two elements of the set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$ are transcendental;
(iii) if Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, then $\gamma^{\left(n\right)}$ is transcendental infinitely often; and
(iv) if Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, then the density of the set of transcendental $\gamma^{\left(n\right)}$ evaluated at $N\in\mathbb{Z}_{\geq 2}$ (i.e., $\left\{n\leq N:\gamma^{\left(n\right)}\textrm{ is transcendental}\right\}/N$) has asymptotic lower bound $\Omega\left(\ln\left(N+2\right)/N\right)$ as $N\rightarrow\infty$.
?proofname?.
(i) Since $\zeta\left(2\right)=\pi^{2}/6$ is transcendental, one can see from the first two lines of (13) that if either $\gamma$ or $\gamma^{\left(2\right)}=\gamma^{2}+\zeta\left(2\right)$ is algebraic, then the other member of the pair must be transcendental.
(ii) We proceed by induction on $m\in\left\{2,3,\ldots,n\right\}$. For the base case, $m=2$, it is known from part (i) that $\mathcal{G}_{2}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}=\left\{\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\gamma^{\left(4\right)}\right\}$ contains at least one transcendental element. For any $m\geq 2$, we show that if $\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}$ contains at least one transcendental element, then it must contain at least two transcendental elements, which in turn implies that $\mathcal{G}_{m+1}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}=\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\gamma^{\left(m\right)}\right\}\right)\cup\left\{\gamma^{\left(2m+1\right)},\gamma^{\left(2m+2\right)}\right\}$ contains at least one transcendental element.
To that end, assume it is known that the set $\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}$ contains at least one transcendental element, and suppose (for purposes of contradiction) that it contains only one such element, denoted by $\theta$; that is, all other elements of $\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}$ are algebraic. It then follows from (17) and Conjecture 1 (via Lemma 5) that there exists a non-trivial polynomial equation in $\theta$ of the form given by (19) with algebraic coefficient $A_{d}^{\left(m,\theta\right)}\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right)$ that does not vanish at $\boldsymbol{\gamma^{*}}$ for at least one $d\geq 1$. Consequently, any root $\theta$ of the relevant equation must be algebraic, contradicting the assumption that only $\theta$ is transcendental and implying that $\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}$ contains at least two transcendental elements. (It is important to note that this approach cannot be extended to show that “at least three” elements of a set of distinct generalized Euler-Mascheroni constants are transcendental because the presence of a second transcendental constant removes the necessity that, for a given $\theta$, all coefficients of $\widetilde{P}_{m,\theta}\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)=0$ are algebraic.)
(iii) Assuming Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, part (ii) implies that the set $\mathcal{G}_{n}^{\left(\textrm{U}\right)}$ contains at least one transcendental element for all $n$. Since $n$ can be arbitrarily large, this precludes the possibility that there exists a finite $n^{*}\in\mathbb{Z}_{\geq 1}$ such that $\gamma^{\left(n\right)}$ is algebraic for all $n\geq n^{*}$.
(iv) Assuming Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, we can identify a lower bound on the frequency with which transcendental values $\gamma^{\left(n\right)}$ can arise by constructing the sparsest feasible subsequence of transcendental elements $\left\{\gamma^{\left(\ell_{i}\right)}\right\}_{i\geq 1}$ satisfying parts (i) and (ii) as follows:
$\bullet\>$ Let $\ell_{1}=1$ and $\ell_{2}=4$.
$\bullet\>$ Given a set of previous element indices $\left\{\ell_{1},\ell_{2},\ldots,\ell_{i-1}\right\}$, select the new element index, $\ell_{i}$, to be as large as possible while still including $\ell_{i-1}+1$ in the interval $\left[\left\lceil\ell_{i}/2\right\rceil,\ell_{i}\right]$ (i.e., $\ell_{i}=2\ell_{i-1}+2$).
This procedure generates the sequence $\left\{\ell_{i}\right\}_{i\geq 1}=\left\{1,4,10,\ldots,3\cdot 2^{i-1}-2,\ldots\right\}$, implying that out of the first $3\cdot 2^{i-1}-2$ positive integers there are at least $i$ element indices – or equivalently, out of the first $N$ positive integers there are at least $\left\lfloor\log_{2}\left(\left(N+2\right)/3\right)+1\right\rfloor=\left\lfloor\ln\left(\left(N+2\right)/3\right)/\ln\left(2\right)+1\right\rfloor$ transcendental elements. As $N\rightarrow\infty$, an asymptotic lower bound on the density of the set of transcendental $\gamma^{\left(n\right)}$ is thus
$\Omega\left(\dfrac{1}{N}\left[\dfrac{\ln\left(N+2\right)-\ln\left(3\right)}{\ln\left(2\right)}+1\right]\right)=\Omega\left(\dfrac{\ln\left(N+2\right)}{N}\right).$
From the $\textrm{Gumbel}\left(0,1\right)$ moment-generating function, $\textrm{M}_{X}\left(t\right)=\Gamma\left(1-t\right),\>t<1$, it is easy to see that
$\gamma^{\left(n\right)}=\textrm{E}_{X}\left[X^{n}\right]=\left(-1\right)^{n}\Gamma^{\left(n\right)}\left(1\right),$
so that all four parts of Theorem 1 are equally valid if one replaces $\gamma^{\left(\bullet\right)}$ with $\Gamma^{\left(\bullet\right)}\left(1\right)$ everywhere. In particular, if Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, then $\Gamma^{\left(n\right)}\left(1\right)$ is transcendental infinitely often, a result consistent with the common expectation that $\Gamma^{\left(n\right)}\left(1\right)$ is transcendental for all $n\geq 1$ (see, e.g., Rivoal [10] and Fischler and Rivoal [11]). [^9] It may be of interest to some readers to note that the asymptotic lower bound on the density of the set of posited transcendental $\Gamma^{\left(n\right)}\left(1\right)$ is smaller (in magnitude) than the asymptotic lower bound on the density of the set of irrational odd-$n$ $\zeta\left(n\right)$ given by Fischler, Sprang, and Zudilin [12]: $\Omega\left(\dfrac{2^{\left(1-\varepsilon\right)\ln\left(N\right)/\ln\left(\ln\left(N\right)\right)}}{N/2}\right)=\Omega\left(N^{\left(1-\varepsilon\right)\ln\left(2\right)/\ln\left(\ln\left(N\right)\right)-1}\right),$ for arbitrarily small $\varepsilon>0$.
3. Algebraic Independence of the $\eta^{\left(n\right)}$
Although the generalized Eta constants do not satisfy simple recurrences such as those derived from (9) and (10), they actually admit more powerful methods for the study of their arithmetic properties. Specifically, we can embed $\eta^{\left(n\right)}=-n!{\textstyle\sum_{k=1}^{\infty}}{\displaystyle\left(-1\right)^{k}/\left(k^{n}\cdot k!\right)}$ into the sequence of functions
$F_{n}\left(z\right)=-n!{\displaystyle\sum_{k=1}^{\infty}}\dfrac{z^{k}}{k^{n}\cdot k!},\quad n\in\mathbb{Z}_{\geq 0},z\in\mathbb{C},$ (22)
which are essentially constant multiples of the polylogarithm-exponential series first studied by Hardy [13] and later used extensively by Shidlovskii [14] in his analysis of $E$-functions. For the purposes at hand, we note that these functions are readily matched to the system $w_{n}\left(z\right)=1+{\textstyle\sum_{k=1}^{\infty}}z^{k}/\left(k^{n}\cdot k!\right)$ of Shidlovskii [14] (Chapter 7, Section 1) via the simple linear identity
$F_{n}\left(z\right)\equiv n!\left(1-w_{n}\left(z\right)\right).$
This allows us to make the following assertions:
(I) The $F_{n}\left(z\right)$, like the $w_{n}\left(z\right)$, are $E$-functions (in the sense of Siegel) that satisfy an $\left(n+1\right)$-order non-homogeneous linear differential equation with coefficients in $\mathbb{Q}\left(z\right)$ and unique finite singularity at $z=0$ (see Chapter 7, Section 1 of Shidlovskii [14]).
(II) For any $m\in\mathbb{Z}_{\geq 1}$ and $z\in\overline{\mathbb{Q}}^{\times}$, the set of values $\left\{F_{0}\left(z\right),F_{1}\left(z\right),\ldots,F_{m}\left(z\right)\right\}$, like the set of values $\left\{w_{0}\left(z\right),w_{1}\left(z\right),\ldots,w_{m}\left(z\right)\right\}$, are algebraically independent over $\overline{\mathbb{Q}}$ (see Chapter 7, Section 1, Theorem 1 of Shidlovskii [14]; the theorem is restated and proved in Chapter 8, Section 3 of the same volume). [^10] These assertions do not explicitly mention that the functions $F_{0}\left(z\right),F_{1}\left(z\right),\ldots,F_{m}\left(z\right)$ are themselves algebraically independent over $\overline{\mathbb{Q}}\left(z\right)$ – a common step in the conventional Siegel-Shidlovskii framework (see, e.g., Beukers [15]). However, the present formulation is faithful to the indicated theorem of Shidlovskii [14], which is tailored for the specific family of $E$-functions considered.
As an immediate consequence of assertions (I) and (II), we obtain the following theorem.
Theorem 2.
The set of values $\left\{\eta^{\left(n\right)}\right\}_{n\geq 0}$ are algebraically independent over $\overline{\mathbb{Q}}$, implying that $\eta^{\left(n\right)}$ is transcendental for all $n\in\mathbb{Z}_{\geq 0}$.
?proofname?.
This result follows directly from the above-mentioned Theorem 1 of Shidlovskii [14]. First, set z=1z=-1 in assertion (II) above to show that, for any $m\in\mathbb{Z}_{\geq 1}$, the set of values $\left\{\eta^{\left(0\right)},\eta^{\left(1\right)},\ldots,\eta^{\left(m\right)}\right\}$ are algebraically independent over $\overline{\mathbb{Q}}$. Next, note that the infinite set $\left\{\eta^{\left(n\right)}\right\}_{n\geq 0}$ are algebraically independent if and only if the subsets $\left\{\eta^{\left(0\right)},\eta^{\left(1\right)},\ldots,\eta^{\left(m\right)}\right\}$ are algebraically independent for all finite $m$. ∎
As indicated above, Theorem 2 is a direct consequence of Shidlovskii’s theorem. However, we have been unable to find an application of this theorem to the generalized partial $\textrm{Gumbel}\left(0,1\right)$ moments of (7) in the literature apart from the specific case of $n=1$ mentioned in Footnote 4. The following corollary shows the impact of the transcendence of the generalized Eta constants on the corresponding generalized Euler-Mascheroni and Euler-Gompertz constants.
Corollary 1.
For all $n\in\mathbb{Z}_{\geq 1}$:
(i) at least one element of the pair $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e\right\}$ is transcendental;
(ii) at least one element of the pair $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}\right\}$ is transcendental; and
(iii) at least two elements of the triple $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e,\delta^{\left(n\right)}\right\}$ are transcendental.
?proofname?.
(i) Theorem 2 states that $\eta^{\left(n\right)}$ is transcendental for all $n\in\mathbb{Z}_{\geq 0}$. Therefore, it is clear from (8) that the two numbers $\gamma^{\left(n\right)}$ and $\delta^{\left(n\right)}/e$ cannot both be algebraic. (We omit the case of $n=0$ from the statement of part (i) because it is trivially obvious that $\delta^{\left(0\right)}/e=-1/e$ is transcendental.)
(ii) From Theorem 2, we know that $\eta^{\left(n\right)}=\gamma^{\left(n\right)}+\delta^{\left(n\right)}/e$ is algebraically independent of $\eta^{\left(0\right)}=1-1/e$ for all $n\in\mathbb{Z}_{\geq 1}$. Therefore, if we assume (for purposes of contradiction) that both $\gamma^{\left(n\right)}$ and $\delta^{\left(n\right)}$ are algebraic, then we can write
$\eta^{\left(n\right)}=\gamma^{\left(n\right)}+\delta^{\left(n\right)}\left(1-\eta^{\left(0\right)}\right)$
$\Longleftrightarrow\eta^{\left(n\right)}-\left(\gamma^{\left(n\right)}+\delta^{\left(n\right)}\right)+\delta^{\left(n\right)}\eta^{\left(0\right)}=0$
$\Longleftrightarrow a\eta^{\left(n\right)}+b\eta^{\left(0\right)}+c=0,$
where $a,b,c\in\overline{\mathbb{Q}}$. This contradicts the algebraic independence of $\eta^{\left(n\right)}$ and $\eta^{\left(0\right)}$, forcing at least one of $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}\right\}$ to be transcendental. As noted in the Introduction, this result was proved by Rivoal [4] for the case of $n=1$.
(iii) Consider the pair $\left\{\delta^{\left(n\right)}/e,\delta^{\left(n\right)}\right\}$. Since $e$ is transcendental, one can see that if either $\delta^{\left(n\right)}/e$ or $\delta^{\left(n\right)}$ is algebraic, then the other member of the pair must be transcendental. Combining this disjunctive transcendence of $\left\{\delta^{\left(n\right)}/e,\delta^{\left(n\right)}\right\}$ with the disjunctive transcendence of both pairs $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e\right\}$ and $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}\right\}$ (from parts (i) and (ii), respectively), it is easy to see that no two elements of the triple $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e,\delta^{\left(n\right)}\right\}$ can both be algebraic. ∎
For $n=1$, the disjunctive transcendence of $\left(\gamma,\delta/e\right)$ complements the disjunctive transcendence of $\left\{\gamma,\delta\right\}$ proved by Rivoal [4]. The new result also draws attention to the fact that, under the probabilistic interpretation of Hardy’s equation, the constant $\delta/e=\textrm{E}_{X}\left[X^{-}\right]=\textrm{E}_{X}\left[\max\left\{-X,0\right\}\right]$ (for $X\sim\textrm{Gumbel}\left(0,1\right)$) is a more natural companion of $\gamma=\textrm{E}_{X}\left[X\right]$ than is $\delta=-\textrm{E}_{X}\left[X\mid X\leq 0\right]$. In Powers [16], we investigated the analytic relationship between $\gamma$ and $\delta$ in (2) by considering linear combinations of the form $\gamma+\alpha\delta$ for $\alpha\in\mathbb{R}^{\times}$. Letting
$S_{\gamma}\coloneqq{\displaystyle\sum_{k=1}^{\infty}\dfrac{\left(-1\right)^{k}\left(!k\right)}{k}}$
and
$S_{\delta}\coloneqq{\displaystyle\sum_{k=1}^{\infty}\left(-1\right)^{k-1}\left(k-1\right)!}$
denote canonical Borel-summable divergent series for $\gamma$ and $\delta$, [^11] The divergent series $S_{\delta}$ is well-known for having been included in the first letter of Srinivasa Ramanujan to G. H. Hardy (see Berndt and Rankin [17]). respectively (where $k=k!{\textstyle\sum_{\ell=0}^{k}}\left(-1\right)^{\ell}/\ell!$ denotes the $k^{\textrm{th}}$ derangement number), it was found that $\alpha=1/e$ is the unique coefficient such that the series $S_{\gamma}+\alpha S_{\delta}$ converges conventionally. This is because the Borel-transform kernels of both $S_{\gamma}$ and $S_{\delta}$ are characterized by unique logarithmic singularities at 1-1 with associated Stokes constants (for the logarithmic-coefficient normalization at the singularity) of 1/e-1/e and $1$, respectively, thus forcing the divergent terms to cancel when combined as $S_{\gamma}+S_{\delta}/e$.
The following theorem relies on both our earlier analysis of the algebraic structure of the polynomials $P_{n}$ (see Section 2.2.1) and the algebraic-independence result of Theorem 2.
Theorem 3.
For the sequence $\left\{\delta^{\left(n\right)}/e\right\}_{n\geq 1}$:
(i) $\delta^{\left(n\right)}/e$ is transcendental infinitely often; and
(ii) the density of the set of transcendental $\delta^{\left(n\right)}/e$ (i.e., $\left\{n\leq N:\delta^{\left(n\right)}/e\textrm{ is transcendental}\right\}/N$) has asymptotic lower bound $1/2$ as $N\rightarrow\infty$.
?proofname?.
(i) Define $\mathcal{G}_{n}=\left\{\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\ldots,\gamma^{\left(n\right)}\right\}$, $\mathcal{H}_{n}=\left\{\eta,\eta^{\left(2\right)},\eta^{\left(3\right)},\ldots,\eta^{\left(n\right)}\right\}$, and $\mathcal{D}_{n}=\left\{\delta/e,\delta^{\left(2\right)}/e,\delta^{\left(3\right)}/e,\ldots,\delta^{\left(n\right)}/e\right\}$ for $n\in\mathbb{Z}_{\geq 1}$, and let $\mathcal{F}_{n}^{\left(\mathcal{G}\right)}=\mathbb{Q}\left(\mathcal{G}_{n}\right)$, $\mathcal{F}_{n}^{\left(\mathcal{H}\right)}=\mathbb{Q}\left(\mathcal{H}_{n}\right)$, and $\mathcal{F}_{n}^{\left(\mathcal{D}\right)}=\mathbb{Q}\left(\mathcal{D}_{n}\right)$ denote the respective fields generated by these three sets of numbers over $\mathbb{Q}$. We wish to demonstrate that
$\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{G}\right)}\right)\leq n-\lfloor n/2\rfloor+1$ (23)
for all $n$, and begin by noting that this inequality holds for $n<4$ because $\mathcal{F}_{n}^{\left(\mathcal{G}\right)}$ is generated by at most $|\mathcal{G}_{n}|$ elements, and $|\mathcal{G}_{n}|\leq n\leq n-\lfloor n/2\rfloor+1$ for $n\in\left\{1,2,3\right\}$. In the following analysis, we treat the case of $n\geq 4$.
From (16), one can see that for any $m\in\left\{2,3,\ldots,n\right\}$, $P_{m}$ is linear in the highest-index constant, $\gamma^{\left(2m\right)}$, with a non-zero integer coefficient $\alpha_{m,0}$. This implies that each $\gamma^{\left(2m\right)}$ can be expressed as a rational polynomial in strictly lower-index constants; that is, $\gamma^{\left(2m\right)}\in\mathbb{Q}\left(\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\ldots,\gamma^{\left(2m-1\right)}\right)$. We now let
$\mathcal{B}_{n}=\mathcal{G}_{n}\setminus\left\{\gamma^{\left(2m\right)}:2\leq m\leq\lfloor n/2\rfloor\right\}$
and show by induction on $m$ that $\gamma^{\left(2m\right)}\in\mathbb{Q}\left(\mathcal{B}_{n}\right)$ for all $m\in\left\{2,3,\ldots,\lfloor n/2\rfloor\right\}$.
For $m=2$, the relation $P_{2}=0$ expresses $\gamma^{\left(4\right)}$ as a rational polynomial in $\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)}\in\mathcal{B}_{n}$, implying $\gamma^{\left(4\right)}\in\mathbb{Q}\left(\mathcal{B}_{n}\right)$. Now consider $m\geq 3$, and assume that $\gamma^{\left(4\right)},\gamma^{\left(6\right)},\ldots,\gamma^{\left(2m-2\right)}\in\mathbb{Q}\left(\mathcal{B}_{n}\right)$. Since (a) $P_{m}=0$ expresses $\gamma^{\left(2m\right)}$ as a rational polynomial in $\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\ldots,\gamma^{\left(2m-1\right)}$, (b) every odd-indexed element of the set $\left\{\gamma^{\left(3\right)},\gamma^{\left(5\right)},\ldots,\gamma^{\left(2m-1\right)}\right\}$ lies in $\mathcal{B}_{n}$, and (c) every even-indexed element of $\left\{\gamma^{\left(4\right)},\gamma^{\left(6\right)},\ldots,\gamma^{\left(2m-2\right)}\right\}$ lies in $\mathbb{Q}\left(\mathcal{B}_{n}\right)$, it follows that $\gamma^{\left(2m\right)}\in\mathbb{Q}\left(\mathcal{B}_{n}\right)$ as well. This confirms that $\gamma^{\left(2m\right)}\in\mathbb{Q}\left(\mathcal{B}_{n}\right)$ for all $m\in\left\{2,3,\ldots,\lfloor n/2\rfloor\right\}$, and thus that $\mathcal{F}_{n}^{\left(\mathcal{G}\right)}=\mathbb{Q}\left(\mathcal{G}_{n}\right)=\mathbb{Q}\left(\mathcal{B}_{n}\right)$.
Given that $\mathcal{B}_{n}$ contains $n-\lfloor n/2\rfloor+1$ elements, the transcendence degree of $\mathcal{F}_{n}^{\left(\mathcal{G}\right)}$ (which is generated by $\mathcal{B}_{n}$) satisfies (23). Then, from the relation in (8), one can see that $\mathcal{F}_{n}^{\left(\mathcal{H}\right)}$ is contained in the composite field $\mathcal{F}_{n}^{\left(\mathcal{G}\right)}\left(\mathcal{F}_{n}^{\left(\mathcal{D}\right)}\right)$, implying
$\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{H}\right)}\right)\leq\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{G}\right)}\left(\mathcal{F}_{n}^{\left(\mathcal{D}\right)}\right)\right)\leq\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{G}\right)}\right)+\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{D}\right)}\right).$ (24)
Since the values in $\mathcal{H}_{n}$ are algebraically independent by Theorem 2, we know that $\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{H}\right)}\right)=n$, which (in conjunction with (23) and (24)) allows us to write
$n\leq n-\lfloor n/2\rfloor+1+\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{D}\right)}\right)$
$\Longleftrightarrow\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{D}\right)}\right)\geq\lfloor n/2\rfloor-1.$
Finally, since the transcendence degree of the field generated by a set of numbers provides a lower bound on the number of transcendental elements in the indicated set, and $\underset{n\rightarrow\infty}{\lim}\left(\lfloor n/2\rfloor-1\right)=\infty$, it follows that the sequence $\left\{\delta^{\left(n\right)}/e\right\}_{n\geq 1}$ must contain an infinite number of transcendental terms.
(ii) The density of the set of transcendental $\delta^{\left(n\right)}/e$ evaluated at $N\in\mathbb{Z}_{\geq 2}$ is given by
$\dfrac{\#\left\{n\leq N:\delta^{\left(n\right)}/e\textrm{ is transcendental}\right\}}{N}\geq\frac{\mathrm{trdeg}\left(\mathcal{F}_{N}^{\left(\mathcal{D}\right)}\right)}{N}\geq\dfrac{\lfloor N/2\rfloor-1}{N}.$
Consequently, as $N\rightarrow\infty$, we have the following asymptotic lower bound on the density of the set of transcendental $\delta^{\left(n\right)}/e$:
$\underset{N\rightarrow\infty}{\liminf}\left(\dfrac{\lfloor N/2\rfloor-1}{N}\right)=\dfrac{1}{2}.$
4. Non-Alternating Analogues
4.1. Two Additional Sequences
Replacing $\left(-1\right)^{k}$ by $\left(1\right)^{k}=1$ in the numerator of the series in (7) gives the “non-alternating analogue” of the sequence $\eta^{\left(n\right)}$,
$\widetilde{\eta}^{\left(n\right)}=-n!{\displaystyle\sum_{k=1}^{\infty}\dfrac{1}{k^{n}\cdot k!}}.$ (25)
We then define the corresponding analogue of the sequence $\delta^{\left(n\right)}$ implicitly through the following natural counterpart to (8),
$\gamma^{\left(n\right)}+\dfrac{\widetilde{\delta}^{\left(n\right)}}{e}=\widetilde{\eta}^{\left(n\right)},$ (26)
where the $\widetilde{\delta}^{\left(n\right)}$ and $\widetilde{\eta}^{\left(n\right)}$ will be called the generalized non-alternating Euler-Gompertz and non-alternating Eta constants, respectively (with $\widetilde{\delta}^{\left(1\right)}=\widetilde{\delta}=-e\textrm{Ei}\left(1\right)$ and $\widetilde{\eta}^{\left(1\right)}=\widetilde{\eta}=\textrm{Ein}\left(-1\right)$ denoting the ordinary non-alternating Euler-Gompertz and non-alternating Eta constants). As a counterpart of (2), we thus have
$\gamma+\dfrac{\widetilde{\delta}}{e}=-{\displaystyle\sum_{k=1}^{\infty}\dfrac{1}{k\cdot k!}}=\textrm{Ein}\left(-1\right),$ (27)
which is shown to be transcendental by Theorem 4 of Section 4.3 below.
4.2. Asymptotic Behavior
Table 2 provides the first 16 values of the generalized non-alternating Euler-Gompertz and non-alternating Eta constants. From its two columns, we can see that the (consistently negative) $\widetilde{\eta}^{\left(n\right)}$ values are comparable in absolute magnitude to the corresponding $\eta^{\left(n\right)}$ of Table 1, and thus increase approximately factorially. However, the (consistently negative) $\widetilde{\delta}^{\left(n\right)}$ values are quite different in magnitude from the $\delta^{\left(n\right)}$ of Table 1 because they, like the $\widetilde{\eta}^{\left(n\right)}$, grow approximately factorially as well.
Table 2. Values of $\widetilde{\delta}^{\left(n\right)}$ and $\widetilde{\eta}^{\left(n\right)}$ for $n\in\left\{0,1,\ldots,15\right\}$
$\boldsymbol{n}$ $\boldsymbol{\widetilde{\delta}^{\left(n\right)}}$ $\boldsymbol{\widetilde{\eta}^{\left(n\right)}}$
0 7.3890560989-7.3890560989 1.7182818285-1.7182818285
11 5.1514643230-5.1514643230 1.3179021515-1.3179021515
22 11.6100810693-11.6100810693 2.2929981451-2.2929981451
33 32.2422478187-32.2422478187 6.4163856532-6.4163856532
44 131.4700021171-131.4700021171 24.8036368256-24.8036368256
55 651.8492520020-651.8492520020 121.9625302861-121.9625302861
66 3,916.6763381264-3,916.6763381264 725.7973399920-725.7973399920
77 27,400.1009939266-27,400.1009939266 5,060.0849690569-5,060.0849690569
88 219,211.5495238585-219,211.5495238585 40,399.8007638273-40,399.8007638273
99 1,972,830.5386794810-1,972,830.5386794810 363,237.5069807080-363,237.5069807080
$10$ 19,728,269.2785592814-19,728,269.2785592814 3,630,582.2647192273-3,630,582.2647192273
$11$ 217,010,407.4543390336-217,010,407.4543390336 39,926,583.2712579089-39,926,583.2712579089
$12$ 2,604,123,546.6077724787-2,604,123,546.6077724787 479,060,223.3022788290-479,060,223.3022788290
$13$ 33,853,598,434.1585762781-33,853,598,434.1585762781 6,227,401,522.0546076015-6,227,401,522.0546076015
$14$ 473,950,346,279.9399036234-473,950,346,279.9399036234 87,180,954,721.7674533138-87,180,954,721.7674533138
$15$ 7,109,255,026,416.1913290094-7,109,255,026,416.1913290094 1,307,694,336,767.4617097988-1,307,694,336,767.4617097988
The following proposition quantifies the asymptotic behavior of both sequences.
Proposition 2.
As $n\rightarrow\infty$:
(i) $\widetilde{\delta}^{\left(n\right)}=-n!\left(2e+e/3^{n+1}+O\left(1/5^{n}\right)\right)$; and
(ii) $\widetilde{\eta}^{\left(n\right)}=-n!\left(1+1/2^{n+1}+O\left(1/3^{n}\right)\right)$.
?proofname?.
See the Appendix. ∎
4.3. Transcendence Results
Despite the clear resemblance of (26) to (8), the former relation does not admit a moment-decomposition interpretation analogous to (4) because the “positive partial moment” component, $\widetilde{\eta}^{\left(n\right)}$, is strictly negative. Moreover, for the case of $n=1$, there does not appear to be anything particularly salient about the coefficient $1/e$ in the linear combination (27) as there was for this coefficient in (2). Thus, the similarity of (27) to (2) arises primarily from the status of $\widetilde{\eta}^{\left(n\right)}$ as a non-alternating analogue of $\eta^{\left(n\right)}$, with $\widetilde{\delta}^{\left(n\right)}$ emerging implicitly by constructing (26) to mimic (8).
Nevertheless, it is quite straightforward to obtain results analogous to Theorem 2, Corollary 1, and Theorem 3, as shown below.
Theorem 4.
The set of values $\left\{\widetilde{\eta}^{\left(n\right)}\right\}_{n\geq 0}$ are algebraically independent over $\overline{\mathbb{Q}}$, implying that $\widetilde{\eta}^{\left(n\right)}$ is transcendental for all $n\in\mathbb{Z}_{\geq 0}$.
?proofname?.
The proof of this result is essentially the same as that of Theorem 2 after setting $z=1$ (rather than z=1z=-1) in assertion (II). [^12] For $n=1$, the present theorem proves the transcendence of $\gamma+\widetilde{\delta}/e=\textrm{Ein}\left(-1\right)$, as noted previously. For this same value of $n$, one can prove the transcendence of $\textrm{Ein}\left(z\right)$ for any $z\in\overline{\mathbb{Q}}^{\times}$ simply by replacing $z$ by z-z in assertion (II).
Corollary 2.
For all $n\in\mathbb{Z}_{\geq 1}$:
(i) at least one element of the pair $\left\{\gamma^{\left(n\right)},\widetilde{\delta}^{\left(n\right)}/e\right\}$ is transcendental;
(ii) at least one element of the pair $\left\{\gamma^{\left(n\right)},\widetilde{\delta}^{\left(n\right)}\right\}$ is transcendental; and
(iii) at least two elements of the triple $\left\{\gamma^{\left(n\right)},\widetilde{\delta}^{\left(n\right)}/e,\widetilde{\delta}^{\left(n\right)}\right\}$ are transcendental.
?proofname?.
The proofs of all three parts of this corollary are entirely analogous to those of the corresponding parts of Corollary 1. ∎
Theorem 5.
For the sequence $\left\{\widetilde{\delta}^{\left(n\right)}/e\right\}_{n\geq 1}$:
(i) $\widetilde{\delta}^{\left(n\right)}/e$ is transcendental infinitely often; and
(ii) the density of the set of transcendental $\widetilde{\delta}^{\left(n\right)}/e$ (i.e., $\left\{n\leq N:\widetilde{\delta}^{\left(n\right)}/e\textrm{ is transcendental}\right\}/N$) has asymptotic lower bound $1/2$ as $N\rightarrow\infty$.
?proofname?.
The proofs of both parts of this theorem are entirely analogous to those of the corresponding parts of Theorem 3. ∎
5. Conclusion
In the present article, we defined sequences of generalized Euler-Mascheroni, Euler-Gompertz, and Eta constants, denoted by $\gamma^{\left(n\right)}$, $\delta^{\left(n\right)}$, and $\eta^{\left(n\right)}$, respectively. After characterizing the basic asymptotic behavior of these sequences, we provided the following results:
$\bullet\>$ at least one element of the pair $\left\{\gamma,\gamma^{\left(2\right)}\right\}$ and – conditional on a realistic conjecture computationally verified for $2\leq n\leq 26$ – at least two elements of each set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$ are transcendental, implying that the $\gamma^{\left(n\right)}$ are transcendental infinitely often, with analogous results holding for the sequence $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$;
$\bullet\>$ the $\eta^{\left(n\right)}$ are algebraically independent, and therefore transcendental, for all $n\geq 0$;
$\bullet\>$ at least one element of each pair, $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e\right\}$ and $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}\right\}$, and at least two elements of the triple $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e,\delta^{\left(n\right)}\right\}$, are transcendental for all $n\geq 1$; and
$\bullet\>$ the $\delta^{\left(n\right)}/e$ are transcendental infinitely often, with the density of the set of transcendental terms having asymptotic lower bound $1/2$.
Subsequently, for the generalized non-alternating Eta and Euler-Gompertz constants, denoted by $\widetilde{\eta}^{\left(n\right)}$ and $\widetilde{\delta}^{\left(n\right)}$, respectively, we found:
$\bullet\>$ the $\widetilde{\eta}^{\left(n\right)}$ are algebraically independent, and therefore transcendental, for all $n\geq 0$;
$\bullet\>$ at least one element of each pair, $\left\{\gamma^{\left(n\right)},\widetilde{\delta}^{\left(n\right)}/e\right\}$ and $\left\{\gamma^{\left(n\right)},\widetilde{\delta}^{\left(n\right)}\right\}$, and at least two elements of the triple $\left\{\gamma^{\left(n\right)},\widetilde{\delta}^{\left(n\right)}/e,\widetilde{\delta}^{\left(n\right)}\right\}$, are transcendental for all $n\geq 1$; and
$\bullet\>$ the $\widetilde{\delta}^{\left(n\right)}/e$ are transcendental infinitely often, with the density of the set of transcendental terms having asymptotic lower bound $1/2$.
Although our primary focus has been to investigate properties of the sequences $\gamma^{\left(n\right)}$, $\delta^{\left(n\right)}$, $\eta^{\left(n\right)}$, $\widetilde{\delta}^{\left(n\right)}$, and $\widetilde{\eta}^{\left(n\right)}$, it is important to note that further insights may be obtained by decomposing the sequence of $E$-functions in (22) by the method of multisections. For example, $F_{n}\left(z\right)$ can be split into the mod-2 multisection components
$G_{n}\left(z\right)=-n!{\displaystyle\sum_{k=1,3,\ldots}\dfrac{z^{k}}{k^{n}\cdot k!}}$
and
$H_{n}\left(z\right)=-n!{\displaystyle\sum_{k=2,4,\ldots}\dfrac{z^{k}}{k^{n}\cdot k!}},$
each of which is itself an $E$-function, and these new functions arranged into a system of linear differential equations with coefficients in $\mathbb{Q}\left(z\right)$ and a unique finite singularity at $z=0$. By employing the Kolchin-Ostrowski differential-field framework (see Srinivasan [18]), we then can show that, for any $n\in\mathbb{Z}_{\geq 1}$, the $G_{0}\left(z\right),H_{0}\left(z\right),G_{1}\left(z\right),H_{1}\left(z\right),\ldots,G_{n}\left(z\right),H_{n}\left(z\right)$ are algebraically independent over $\mathbb{C}\left(z\right)$. This allows us to conclude, via the refined Siegel-Shidlovskii theorem (see Beukers [15]), that for $z=1$ the set of values $\left\{G_{0}\left(1\right),H_{0}\left(1\right),G_{1}\left(1\right),H_{1}\left(1\right),\ldots,G_{n}\left(1\right),H_{n}\left(1\right)\right\}$ are algebraically independent over $\overline{\mathbb{Q}}$, and therefore transcendental. Finally, recognizing that
$G_{n}\left(1\right)=-n!{\displaystyle\sum_{k=1,3,\ldots}\dfrac{1}{k^{n}\cdot k!}=\dfrac{1}{2}\left(\widetilde{\eta}^{\left(n\right)}-\eta^{\left(n\right)}\right)=\dfrac{1}{2}\left(\dfrac{\widetilde{\delta}^{\left(n\right)}}{e}-\dfrac{\delta^{\left(n\right)}}{e}\right)}$
and
$H_{n}\left(1\right)=-n!{\displaystyle\sum_{k=2,4,\ldots}\dfrac{1}{k^{n}\cdot k!}=\dfrac{1}{2}\left(\widetilde{\eta}^{\left(n\right)}+\eta^{\left(n\right)}\right)=\dfrac{1}{2}\left(2\gamma^{\left(n\right)}+\dfrac{\widetilde{\delta}^{\left(n\right)}}{e}+\dfrac{\delta^{\left(n\right)}}{e}\right)},$
we can use the fact that $\widetilde{\delta}^{\left(n\right)}/e-\delta^{\left(n\right)}/e$ is transcendental to demonstrate that at least two elements of the triple $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e,\widetilde{\delta}^{\left(n\right)}/e\right\}$ are transcendental for all $n\geq 1$.
?refname?
  • [1] Aptekarev, A. I., 2009, “On Linear Forms Containing the Euler Constant”, arXiv:0902.1768v2.
  • [2] Shidlovskii, A. B., 1962, “On Transcendency and Algebraic Independence of the Values of $E$-Functions Connected with an Arbitrary Number of Algebraic Equations in the Field of Rational Functions”, Proceedings of the USSR Academy of Sciences, Mathematics Series, 26, 877-910 (in Russian).
  • [3] Mahler, K., 1968, “Applications of a Theorem by A. B. Shidlovski”, Proceedings of the Royal Society A, 305, 1481, 149-173.
  • [4] Rivoal, T., 2012, “On the Arithmetic Nature of the Values of the Gamma Function, Euler’s Constant, and Gompertz’s Constant”, Michigan Mathematics Journal, 61, 2, 239-254.
  • [5] Lagarias, J. C., 2013, “Euler’s Constant: Euler’s Work and Modern Developments”, Bulletin of the American Mathematical Society, 50, 4, 527-628.
  • [6] Hardy, G. H., 1949, Divergent Series, Clarendon Press, Oxford, UK.
  • [7] Powers, M. R., 2025, “Conjunctions of Three ’Euler Constants’ in Poisson-Related Expressions”, arXiv:2503.02054.
  • [8] Pinkus, A., 2010, Totally Positive Matrices, Cambridge University Press, Cambridge, UK.
  • [9] Liu, Q., 2002, Algebraic Geometry and Arithmetic Curves, Oxford University Press, Oxford, UK.
  • [10] Rivoal, T., 2009, “Rational Approximations for Values of Derivatives of the Gamma Function”, Transactions of the American Mathematical Society, 361, 11, 6115-6149.
  • [11] Fischler, S. and Rivoal, T., 2024, “Relations between Values of Arithmetic Gevrey Series, and Applications to Values of the Gamma Function”, Journal of Number Theory, 261, 36-54.
  • [12] Fischler, S., Sprang, J., and Zudilin, W., 2019, “Many Odd Zeta Values are Irrational”, Compositio Mathematica, 155, 938-952.
  • [13] Hardy, G. H., 1905, “On the Zeroes of Certain Classes of Integral Taylor Series, Part II”, Proceedings of the London Mathematical Society, s2-2, 1, 401-431.
  • [14] Shidlovskii, A. B., 1989, Transcendental Numbers, Walter de Gruyter, Berlin, Germany.
  • [15] Beukers, F., 2006, “A Refined Version of the Siegel-Shidlovskii Theorem”, Annals of Mathematics, 163, 369-379.
  • [16] Powers, M. R., 2025, “On the Uniqueness of Ein(1) among Linear Combinations of the Euler-Mascheroni and Euler-Gompertz Constants”, arXiv:2510.00315.
  • [17] Berndt, B. C. and Rankin, R. A., 1995, Ramanujan: Letters and Commentary, American Mathematical Society.
  • [18] Srinivasan, V. R., 2008, “Extensions by Antiderivatives, Exponentials of Integrals, and by Iterated Logarithms”, arXiv:0811.3004.
  • [19]
Appendix
Proof of Proposition 1.
We consider the three parts of the proposition in the order: (iii), (ii), (i).
(iii) From (7), it is easy to see that
$\eta^{\left(n\right)}<n!\left[1-\dfrac{1}{2^{n+1}}+\left(\dfrac{1}{6}\right)\dfrac{1}{3^{n}}+\left(\dfrac{1}{24}\right)\dfrac{1}{4^{n}}+\cdots\right]$
$n!\left[1-\dfrac{1}{2^{n+1}}+\left(\dfrac{1}{6}+\dfrac{1}{24}+\cdots\right)\dfrac{1}{3^{n}}\right]$
$n!\left[1-\dfrac{1}{2^{n+1}}+\dfrac{\left(e-5/2\right)}{3^{n}}\right]$
and
$\eta^{\left(n\right)}>n!\left[1-\dfrac{1}{2^{n+1}}-\left(\dfrac{1}{6}\right)\dfrac{1}{3^{n}}-\left(\dfrac{1}{24}\right)\dfrac{1}{4^{n}}-\cdots\right]$
$n!\left[1-\dfrac{1}{2^{n+1}}-\left(\dfrac{1}{6}+\dfrac{1}{24}+\cdots\right)\dfrac{1}{3^{n}}\right]$
$n!\left[1-\dfrac{1}{2^{n+1}}-\dfrac{\left(e-5/2\right)}{3^{n}}\right].$
It then follows that
$\eta^{\left(n\right)}=n!\left(1-\dfrac{1}{2^{n+1}}+O\left(\dfrac{1}{3^{n}}\right)\right).$
(ii) Substituting $u=e^{-x}$ into the integral in (6) yields
$\delta^{\left(n\right)}=\left(-1\right)^{n+1}e{\displaystyle\int_{1}^{\infty}\left[\ln\left(u\right)\right]^{n}e^{-u}du}$
$\Longrightarrow\left|\delta^{\left(n\right)}\right|=e{\displaystyle\int_{1}^{\infty}\left[\ln\left(u\right)\right]^{n}e^{-u}du}=e{\displaystyle\int_{1}^{\infty}e^{\phi_{n}\left(u\right)}du},\qquad\qquad\textrm{(A1)}$
where $\phi_{n}\left(u\right)=n\ln\left(\ln\left(u\right)\right)-u$. We then employ Laplace’s (saddle-point) method to approximate this integral.
Taking derivatives of $\phi_{n}\left(u\right)$ with respect to $u$ gives
$\phi_{n}^{\prime}\left(u\right)=\dfrac{n}{u\ln\left(u\right)}-1\>\textrm{and }\>\phi_{n}^{\prime\prime}\left(u\right)=-\dfrac{n}{u^{2}\ln\left(u\right)}\left(1+\dfrac{1}{\ln\left(u\right)}\right)<0,$
revealing that $\phi_{n}\left(u\right)$ enjoys a unique global maximum at the saddle point
$u^{*}\ln\left(u^{*}\right)=n\Longleftrightarrow u^{*}=\dfrac{n}{W\left(n\right)},\>\ln\left(u^{*}\right)=W\left(n\right).$
Then
$\phi_{n}\left(u^{*}\right)=n\ln\left(\ln\left(u^{*}\right)\right)-\dfrac{n}{\ln\left(u^{*}\right)}=n\ln\left(W\left(n\right)\right)-\dfrac{n}{W\left(n\right)},$
$\left|\phi_{n}^{\prime\prime}\left(u^{*}\right)\right|=\dfrac{W\left(n\right)}{n}\left(1+\dfrac{1}{W\left(n\right)}\right)=\dfrac{W\left(n\right)+1}{n},$
and (A1) can be approximated by
$\left|\delta^{\left(n\right)}\right|=e\cdot e^{\phi_{n}\left(u^{*}\right)}\sqrt{\dfrac{2\pi}{\left|\phi_{n}^{\prime\prime}\left(u^{*}\right)\right|}}\left(1+o\left(1\right)\right)$
$e\left[W\left(n\right)\right]^{n}\exp\left(-\dfrac{n}{W\left(n\right)}\right)\sqrt{\dfrac{2\pi n}{W\left(n\right)+1}}\left(1+o\left(1\right)\right).$
(i) Finally, we assemble the results in parts (iii) and (ii) via (8), giving
$\gamma^{\left(n\right)}=n!\left(1-\dfrac{1}{2^{n}\cdot 2!}+\dfrac{1}{3^{n}\cdot 3!}-\dfrac{1}{4^{n}\cdot 4!}+\cdots\right)-\dfrac{1}{e}\left(-1\right)^{n+1}e{\displaystyle\int_{1}^{\infty}\left[\ln\left(u\right)\right]^{n}e^{-u}du}$
$n!\left(1-\dfrac{1}{2^{n+1}}+O\left(\dfrac{1}{3^{n}}\right)\right)-\left(-1\right)^{n+1}\left[W\left(n\right)\right]^{n}\exp\left(-\dfrac{n}{W\left(n\right)}\right)\sqrt{\dfrac{2\pi n}{W\left(n\right)+1}}\left(1+o\left(1\right)\right)$
$n!\left(1-\dfrac{1}{2^{n+1}}+O\left(\dfrac{1}{3^{n}}\right)\right).$
Proof of Proposition 2.
We consider the two parts of the proposition in reverse order.
(ii) From (25), it follows that
$\widetilde{\eta}^{\left(n\right)}>-n!\left[1+\dfrac{1}{2^{n+1}}+\left(\dfrac{1}{6}+\dfrac{1}{24}+\cdots\right)\dfrac{1}{3^{n}}\right]$
$n!\left[1+\dfrac{1}{2^{n+1}}+\dfrac{\left(e-5/2\right)}{3^{n}}\right]$
and
$\widetilde{\eta}^{\left(n\right)}<-n!\left[1+\dfrac{1}{2^{n+1}}+\left(\dfrac{1}{6}\right)\dfrac{1}{3^{n}}\right],$
implying
$\widetilde{\eta}^{\left(n\right)}=-n!\left(1+\dfrac{1}{2^{n+1}}+O\left(\dfrac{1}{3^{n}}\right)\right).$
(i) Now solve for $\widetilde{\delta}^{\left(n\right)}$ via (26), using the initial expression for $\gamma^{\left(n\right)}$ in the proof of Proposition 1(i):
$\begin{array}[]{c}\widetilde{\delta}^{\left(n\right)}=e\left[-n!\left(1+\dfrac{1}{2^{n}\cdot 2!}+\dfrac{1}{3^{n}\cdot 3!}+\dfrac{1}{4^{n}\cdot 4!}+\cdots\right)-n!\left(1-\dfrac{1}{2^{n}\cdot 2!}+\dfrac{1}{3^{n}\cdot 3!}-\dfrac{1}{4^{n}\cdot 4!}+\cdots\right)\right.\\ \left.+\dfrac{1}{e}\left(-1\right)^{n+1}e{\displaystyle\int_{1}^{\infty}\left[\ln\left(u\right)\right]^{n}e^{-u}du}\right]\end{array}$
$e\left[-2n!\left(1+\dfrac{1}{2\cdot 3^{n+1}}+O\left(\dfrac{1}{5^{n}}\right)\right)\right]$
$n!\left(2e+\dfrac{e}{3^{n+1}}+O\left(\dfrac{1}{5^{n}}\right)\right).$

### v5

@@
 The two constants are related through a well-known equation of Hardy, equivalent to $\gamma+\delta/e=\textrm{Ein}\left(1\right)$, which recently has been generalized to $\gamma^{\left(n\right)}+\delta^{\left(n\right)}/e=\eta^{\left(n\right)},\>n\geq 0$ for sequences of constants $\gamma^{\left(n\right)}$, $\delta^{\left(n\right)}$, and $\eta^{\left(n\right)}$ (given respectively by raw, conditional, and partial moments of the Gumbel(0,1) probability distribution).
-Investigating the $\gamma^{\left(n\right)}$ through recurrence relations (where $\gamma^{\left(0\right)}=1$ and $\gamma^{\left(1\right)}=\gamma$), we find that at least one of the pair $\{\gamma,\gamma^{\left(2\right)}\}$ and – conditional on a realistic conjecture verified for $2\leq n\leq 26$ – at least two of each set $\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\}$ are transcendental, implying that the $\gamma^{\left(n\right)}$ are transcendental infinitely often (with analogous results for the sequence $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$).
+Investigating the $\gamma^{\left(n\right)}$ through recurrence relations (where $\gamma^{\left(0\right)}=1$ and $\gamma^{\left(1\right)}=\gamma$), we find that at least one of the pair $\{\gamma,\gamma^{\left(2\right)}\}$ and at least two of each set $\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\}$, for $2\leq n\leq 26$, are transcendental, with analogous results for the sequence $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$.
 We then show, via a theorem of Shidlovskii, that the $\eta^{\left(n\right)}$ are algebraically independent, and therefore transcendental, for all $n\geq 0$, implying that at least one of each pair, $\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e\}$ and $\{\gamma^{\left(n\right)},\delta^{\left(n\right)}\}$, and at least two of the triple $\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e,\delta^{\left(n\right)}\}$, are transcendental for all $n\geq 1$.
@@
 In Section 2, we begin our study of the arithmetic properties of the three generalized-constant sequences by investigating recurrence relations associated with the $\gamma^{\left(n\right)}$.
-These recurrences reveal that at least one element of the pair $\left\{\gamma,\gamma^{\left(2\right)}\right\}$ and – conditional on a realistic conjecture computationally verified for $2\leq n\leq 26$ – at least two elements of each set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$ are transcendental, implying that the $\gamma^{\left(n\right)}$ are transcendental infinitely often (with analogous results for the sequence $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$).
-Next, in Section 3, we apply a theorem of Shidlovskii (1989) to demonstrate the algebraic independence, and therefore transcendence, of the $\eta^{\left(n\right)}$ for all $n\geq 0$.
+These recurrences reveal that at least one element of the pair $\left\{\gamma,\gamma^{\left(2\right)}\right\}$ and at least two elements of each set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$, for $n\in\left\{2,3,\ldots,26\right\}$, are transcendental.
+As will be seen, the latter result follows from verification, for $n\in\left\{2,3,\ldots,26\right\}$, of a certain reasonable conjecture (which, if true for all $n\in\mathbb{Z}_{\geq 2}$, necessitates that the $\gamma^{\left(n\right)}$ are transcendental infinitely often).
+Since $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$, all conclusions for the $\gamma^{\left(n\right)}$ apply equally well to the sequence $\Gamma^{\left(n\right)}\left(1\right)$.
+In Section 3, we employ a theorem of Shidlovskii (1989) to demonstrate the algebraic independence, and therefore transcendence, of the $\eta^{\left(n\right)}$ for all $n\geq 0$.
 This implies that at least one element of each pair, $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e\right\}$ and $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}\right\}$, and at least two elements of the triple $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e,\delta^{\left(n\right)}\right\}$, are transcendental for all $n\geq 1$.
 Moreover, further analysis of the $\gamma^{\left(n\right)}$ and $\eta^{\left(n\right)}$ shows that the values $\delta^{\left(n\right)}/e$ are transcendental infinitely often, with the density of the set of transcendental terms having asymptotic lower bound $1/2$.
-Finally, in Section 4 we consider two additional sequences, $\widetilde{\delta}^{\left(n\right)}$ and $\widetilde{\eta}^{\left(n\right)}$, satisfying the “non-alternating analogue” equation $\gamma^{\left(n\right)}+\widetilde{\delta}^{\left(n\right)}/e=\widetilde{\eta}^{\left(n\right)}$, and provide results parallel to those for the original system.
+Finally, Section 4 addresses two additional sequences, $\widetilde{\delta}^{\left(n\right)}$ and $\widetilde{\eta}^{\left(n\right)}$, satisfying the “non-alternating analogue” equation $\gamma^{\left(n\right)}+\widetilde{\delta}^{\left(n\right)}/e=\widetilde{\eta}^{\left(n\right)}$, and provides results parallel to those for the original system.
 2. Recurrence Relations for $\gamma^{\left(n\right)}$
@@
 Assume (for purposes of contradiction) that the polynomials $\left\{P_{2},P_{3},\ldots,P_{n}\right\}$ are algebraically dependent; that is, there exists a non-zero polynomial $Q_{n}\in\mathbb{Q}\left[t_{2},t_{3},\ldots,t_{n}\right]$ such that $Q_{n}\left(P_{2},P_{3},\ldots,P_{n}\right)=0$.
+[^8] If $Q_{n}$ does not actually involve the variable $t_{n}$, then $Q_{n}\in\mathbb{Q}\left[t_{2},t_{3},\ldots,t_{n-1}\right]$ and $Q_{n}\left(P_{2},P_{3},\ldots,P_{n}\right)=0$ already embodies an algebraic dependence among $\left\{P_{2},P_{3},\ldots,P_{n-1}\right\}$.
+In that case, we can replace $n$ by n1n-1 (and repeat this reduction if necessary) so that the assumed dependence relation involves the highest-index variable, $n^{*}$.
+Then, without loss of generality, we take $n=n^{*}$.
 Writing $Q_{n}$ as a polynomial in $t_{n}$ gives
@@
 Although we know there exists $\widetilde{P}_{n,\theta}\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)=0$ that can be expressed as a non-trivial polynomial equation in $\theta\in\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)$ of the form given by (19), it remains possible that $A_{d}^{\left(n,\theta\right)}=0$ for all $d\in\left\{1,2,\ldots,D_{n,\theta}\right\}$ at the actual realizations of the elements $\gamma^{\left(k\right)}\in\mathcal{G}_{n}^{\left(2n\right)}$ (to be denoted by $\boldsymbol{\gamma^{*}}\Longleftrightarrow\gamma^{\left(k\right)}=\textrm{E}_{X}\left[X_{\textrm{G}}^{k}\right]\>\forall\>k\in\mathbb{Z}_{\geq 1}$).
-Lemma 5 below, supported by Conjecture 1, ensures this does not happen.
+Lemma 5 below (supported by the subsequent Conjecture 1) ensures this does not happen.
 Recalling $\mathcal{R}_{n}=\mathbb{Q}\left[\mathcal{G}_{n}^{\left(2n\right)}\right]$ and $\mathcal{I}_{n}=\left\langle P_{2},P_{3},\ldots,P_{n}\right\rangle\subset\mathcal{R}_{n}$ from Section 2.2.1, let $\mathcal{S}_{n}=\mathcal{R}_{n}/\mathcal{I}_{n}$ and define $\mathcal{J}_{n,\boldsymbol{\gamma^{*}}}\subset\mathcal{R}_{n}$ as the kernel of the evaluation homomorphism induced by $\boldsymbol{\gamma^{*}}$ on $\mathcal{R}_{n}$.
@@
 Clearly, greater values of $n$ could be tested, but the program’s runtime increases approximately exponentially in $n$.
-(For example, $n=30$ would require a runtime of about 1 week.) In addition to computational verification, the conjecture is realistic on theoretical grounds because it holds for any given $n$ unless $\boldsymbol{\gamma^{*}}$ is forced to lie on one of finitely many special hypersurfaces (i.e., the vanishing loci of the minors $\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)$).
+(For example, $n=30$ would require a runtime of about 1 week.) In addition to its computational verification, the conjecture is realistic on theoretical grounds because it holds for any given $n$ unless $\boldsymbol{\gamma^{*}}$ is forced to lie on one of finitely many special hypersurfaces (i.e., the vanishing loci of the minors $\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)$).
 Conjecture 1.
@@
 (iii) if Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, then $\gamma^{\left(n\right)}$ is transcendental infinitely often; and
-(iv) if Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, then the density of the set of transcendental $\gamma^{\left(n\right)}$ evaluated at $N\in\mathbb{Z}_{\geq 2}$ (i.e., $\left\{n\leq N:\gamma^{\left(n\right)}\textrm{ is transcendental}\right\}/N$) has asymptotic lower bound $\Omega\left(\ln\left(N+2\right)/N\right)$ as $N\rightarrow\infty$.
+(iv) if Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, then the density of the set of transcendental $\gamma^{\left(n\right)}$ evaluated at $N\in\mathbb{Z}_{\geq 2}$ (i.e., $\left\{n\leq N:\gamma^{\left(n\right)}\textrm{ is transcendental}\right\}/N$) has asymptotic lower bound $\Omega\left(\ln\left(N\right)/N\right)$ as $N\rightarrow\infty$.
 ?proofname?.
@@
 Since $n$ can be arbitrarily large, this precludes the possibility that there exists a finite $n^{*}\in\mathbb{Z}_{\geq 1}$ such that $\gamma^{\left(n\right)}$ is algebraic for all $n\geq n^{*}$.
-(iv) Assuming Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, we can identify a lower bound on the frequency with which transcendental values $\gamma^{\left(n\right)}$ can arise by constructing the sparsest feasible subsequence of transcendental elements $\left\{\gamma^{\left(\ell_{i}\right)}\right\}_{i\geq 1}$ satisfying parts (i) and (ii) as follows:
-$\bullet\>$ Let $\ell_{1}=1$ and $\ell_{2}=4$.
-$\bullet\>$ Given a set of previous element indices $\left\{\ell_{1},\ell_{2},\ldots,\ell_{i-1}\right\}$, select the new element index, $\ell_{i}$, to be as large as possible while still including $\ell_{i-1}+1$ in the interval $\left[\left\lceil\ell_{i}/2\right\rceil,\ell_{i}\right]$ (i.e., $\ell_{i}=2\ell_{i-1}+2$).
-This procedure generates the sequence $\left\{\ell_{i}\right\}_{i\geq 1}=\left\{1,4,10,\ldots,3\cdot 2^{i-1}-2,\ldots\right\}$, implying that out of the first $3\cdot 2^{i-1}-2$ positive integers there are at least $i$ element indices – or equivalently, out of the first $N$ positive integers there are at least $\left\lfloor\log_{2}\left(\left(N+2\right)/3\right)+1\right\rfloor=\left\lfloor\ln\left(\left(N+2\right)/3\right)/\ln\left(2\right)+1\right\rfloor$ transcendental elements.
-As $N\rightarrow\infty$, an asymptotic lower bound on the density of the set of transcendental $\gamma^{\left(n\right)}$ is thus
-$\Omega\left(\dfrac{1}{N}\left[\dfrac{\ln\left(N+2\right)-\ln\left(3\right)}{\ln\left(2\right)}+1\right]\right)=\Omega\left(\dfrac{\ln\left(N+2\right)}{N}\right).$
-From the $\textrm{Gumbel}\left(0,1\right)$ moment-generating function, $\textrm{M}_{X}\left(t\right)=\Gamma\left(1-t\right),\>t<1$, it is easy to see that
+(iv) As previously observed, if Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, then part (ii) forces each set $\mathcal{G}_{n}^{\left(\textrm{U}\right)}$ to contain at least one transcendental element $\gamma^{\left(k_{n}\right)},\>k_{n}\in\left\{n,n+1,\ldots,2n\right\}$.
+Taking $m=3^{i}$ for $i\in\mathbb{Z}_{\geq 1}$, we now consider the family of disjoint intervals $\left[3^{i},2\cdot 3^{i}\right]\ni k_{m}$ and note that we can choose $N\in\mathbb{Z}_{\geq 2}$ such that $N\geq 2\cdot 3^{i}=2m$ and the $k_{m}\in\left[3^{i},2\cdot 3^{i}\right]\subseteq\left\{1,2,\ldots,N\right\}$ are distinct.
+This implies that the density of the set of transcendental $\gamma^{\left(n\right)}$ evaluated at $N$ is given by
+$\dfrac{\#\left\{n\leq N:\gamma^{\left(n\right)}\textrm{ is transcendental}\right\}}{N}\geq\dfrac{\#\left\{k_{m}\leq N:\gamma^{\left(k_{m}\right)}\textrm{ is transcendental}\right\}}{N}$
+$\geq\dfrac{1}{N}\left\lfloor\log_{3}\left(\dfrac{N}{2}\right)\right\rfloor\geq\dfrac{1}{N}\left(\dfrac{\ln\left(N\right)-\ln\left(2\right)}{\ln\left(3\right)}-1\right),$
+with asymptotic lower bound $\Omega\left(\ln\left(N\right)/N\right)$ as $N\rightarrow\infty$. ∎
+Since Conjecture 1 has been verified computationally for $n\in\left\{2,3,\ldots,26\right\}$, we know from part (ii) of the above theorem that at least two elements of each associated set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$ are transcendental.
+Furthermore, it is easy to see from the $\textrm{Gumbel}\left(0,1\right)$ moment-generating function, $\textrm{M}_{X}\left(t\right)=\Gamma\left(1-t\right),\>t<1$, that
 $\gamma^{\left(n\right)}=\textrm{E}_{X}\left[X^{n}\right]=\left(-1\right)^{n}\Gamma^{\left(n\right)}\left(1\right),$
-so that all four parts of Theorem 1 are equally valid if one replaces $\gamma^{\left(\bullet\right)}$ with $\Gamma^{\left(\bullet\right)}\left(1\right)$ everywhere.
-In particular, if Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, then $\Gamma^{\left(n\right)}\left(1\right)$ is transcendental infinitely often, a result consistent with the common expectation that $\Gamma^{\left(n\right)}\left(1\right)$ is transcendental for all $n\geq 1$ (see, e.g., Rivoal [10] and Fischler and Rivoal [11]).
+implying that all four parts of Theorem 1 are equally valid if one replaces $\gamma^{\left(\bullet\right)}$ with $\Gamma^{\left(\bullet\right)}\left(1\right)$ everywhere.
+In particular: (a) at least two elements of each set $\left\{\Gamma^{\left(1\right)}\left(1\right),\Gamma^{\left(n\right)}\left(1\right),\Gamma^{\left(n+1\right)}\left(1\right),\ldots,\Gamma^{\left(2n\right)}\left(1\right)\right\}$ are transcendental for $n\in\left\{2,3,\ldots,26\right\}$; and (b) if Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, then $\Gamma^{\left(n\right)}\left(1\right)$ must be transcendental infinitely often, a result consistent with the common expectation that $\Gamma^{\left(n\right)}\left(1\right)$ is transcendental for all $n\geq 1$ (see, e.g., Rivoal [10] and Fischler and Rivoal [11]).
 [^10] It may be of interest to some readers to note that the asymptotic lower bound on the density of the set of posited transcendental $\Gamma^{\left(n\right)}\left(1\right)$ is smaller (in magnitude) than the asymptotic lower bound on the density of the set of irrational odd-$n$ $\zeta\left(n\right)$ given by Fischler, Sprang, and Zudilin [12]: $\Omega\left(\dfrac{2^{\left(1-\varepsilon\right)\ln\left(N\right)/\ln\left(\ln\left(N\right)\right)}}{N/2}\right)=\Omega\left(N^{\left(1-\varepsilon\right)\ln\left(2\right)/\ln\left(\ln\left(N\right)\right)-1}\right),$ for arbitrarily small $\varepsilon>0$.
@@
 From (16), one can see that for any $m\in\left\{2,3,\ldots,n\right\}$, $P_{m}$ is linear in the highest-index constant, $\gamma^{\left(2m\right)}$, with a non-zero integer coefficient $\alpha_{m,0}$.
-This implies that each $\gamma^{\left(2m\right)}$ can be expressed as a rational polynomial in strictly lower-index constants; that is, $\gamma^{\left(2m\right)}\in\mathbb{Q}\left(\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\ldots,\gamma^{\left(2m-1\right)}\right)$.
+This implies that each $\gamma^{\left(2m\right)}$ can be expressed as a polynomial with rational coefficients in strictly lower-index constants; that is, $\gamma^{\left(2m\right)}\in\mathbb{Q}\left(\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\ldots,\gamma^{\left(2m-1\right)}\right)$.
 We now let
@@
 and show by induction on $m$ that $\gamma^{\left(2m\right)}\in\mathbb{Q}\left(\mathcal{B}_{n}\right)$ for all $m\in\left\{2,3,\ldots,\lfloor n/2\rfloor\right\}$.
-For $m=2$, the relation $P_{2}=0$ expresses $\gamma^{\left(4\right)}$ as a rational polynomial in $\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)}\in\mathcal{B}_{n}$, implying $\gamma^{\left(4\right)}\in\mathbb{Q}\left(\mathcal{B}_{n}\right)$.
+For $m=2$, the relation $P_{2}=0$ expresses $\gamma^{\left(4\right)}$ as a polynomial with rational coefficients in $\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)}\in\mathcal{B}_{n}$, implying $\gamma^{\left(4\right)}\in\mathbb{Q}\left(\mathcal{B}_{n}\right)$.
 Now consider $m\geq 3$, and assume that $\gamma^{\left(4\right)},\gamma^{\left(6\right)},\ldots,\gamma^{\left(2m-2\right)}\in\mathbb{Q}\left(\mathcal{B}_{n}\right)$.
-Since (a) $P_{m}=0$ expresses $\gamma^{\left(2m\right)}$ as a rational polynomial in $\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\ldots,\gamma^{\left(2m-1\right)}$, (b) every odd-indexed element of the set $\left\{\gamma^{\left(3\right)},\gamma^{\left(5\right)},\ldots,\gamma^{\left(2m-1\right)}\right\}$ lies in $\mathcal{B}_{n}$, and (c) every even-indexed element of $\left\{\gamma^{\left(4\right)},\gamma^{\left(6\right)},\ldots,\gamma^{\left(2m-2\right)}\right\}$ lies in $\mathbb{Q}\left(\mathcal{B}_{n}\right)$, it follows that $\gamma^{\left(2m\right)}\in\mathbb{Q}\left(\mathcal{B}_{n}\right)$ as well.
+Since (a) $P_{m}=0$ expresses $\gamma^{\left(2m\right)}$ as a polynomial with rational coefficients in $\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\ldots,\gamma^{\left(2m-1\right)}$, (b) every odd-indexed element of the set $\left\{\gamma^{\left(3\right)},\gamma^{\left(5\right)},\ldots,\gamma^{\left(2m-1\right)}\right\}$ lies in $\mathcal{B}_{n}$, and (c) every even-indexed element of $\left\{\gamma^{\left(4\right)},\gamma^{\left(6\right)},\ldots,\gamma^{\left(2m-2\right)}\right\}$ lies in $\mathbb{Q}\left(\mathcal{B}_{n}\right)$, it follows that $\gamma^{\left(2m\right)}\in\mathbb{Q}\left(\mathcal{B}_{n}\right)$ as well.
 This confirms that $\gamma^{\left(2m\right)}\in\mathbb{Q}\left(\mathcal{B}_{n}\right)$ for all $m\in\left\{2,3,\ldots,\lfloor n/2\rfloor\right\}$, and thus that $\mathcal{F}_{n}^{\left(\mathcal{G}\right)}=\mathbb{Q}\left(\mathcal{G}_{n}\right)=\mathbb{Q}\left(\mathcal{B}_{n}\right)$.
@@
 $\dfrac{\#\left\{n\leq N:\delta^{\left(n\right)}/e\textrm{ is transcendental}\right\}}{N}\geq\frac{\mathrm{trdeg}\left(\mathcal{F}_{N}^{\left(\mathcal{D}\right)}\right)}{N}\geq\dfrac{\lfloor N/2\rfloor-1}{N}.$
-Consequently, as $N\rightarrow\infty$, we have the following asymptotic lower bound on the density of the set of transcendental $\delta^{\left(n\right)}/e$:
+Thus, as $N\rightarrow\infty$, the density has asymptotic lower bound
 $\underset{N\rightarrow\infty}{\liminf}\left(\dfrac{\lfloor N/2\rfloor-1}{N}\right)=\dfrac{1}{2}.$
@@
 In the present article, we defined sequences of generalized Euler-Mascheroni, Euler-Gompertz, and Eta constants, denoted by $\gamma^{\left(n\right)}$, $\delta^{\left(n\right)}$, and $\eta^{\left(n\right)}$, respectively.
-After characterizing the basic asymptotic behavior of these sequences, we provided the following results:
-$\bullet\>$ at least one element of the pair $\left\{\gamma,\gamma^{\left(2\right)}\right\}$ and – conditional on a realistic conjecture computationally verified for $2\leq n\leq 26$ – at least two elements of each set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$ are transcendental, implying that the $\gamma^{\left(n\right)}$ are transcendental infinitely often, with analogous results holding for the sequence $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$;
+After characterizing the basic asymptotic behavior of these sequences, we provided the following (unconditional) results:
+$\bullet\>$ at least one element of the pair $\left\{\gamma,\gamma^{\left(2\right)}\right\}$ and at least two elements of each set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$, for $n\in\left\{2,3,\ldots,26\right\}$, are transcendental, with analogous implications for the sequence $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$;
 $\bullet\>$ the $\eta^{\left(n\right)}$ are algebraically independent, and therefore transcendental, for all $n\geq 0$;
@@
 Although our primary focus has been to investigate properties of the sequences $\gamma^{\left(n\right)}$, $\delta^{\left(n\right)}$, $\eta^{\left(n\right)}$, $\widetilde{\delta}^{\left(n\right)}$, and $\widetilde{\eta}^{\left(n\right)}$, it is important to note that further insights may be obtained by decomposing the sequence of $E$-functions in (22) by the method of multisections.
-For example, $F_{n}\left(z\right)$ can be split into the mod-2 multisection components
+For example, $F_{n}\left(z\right),\>n\in\mathbb{Z}_{\geq 0}$ can be split into the mod-2 multisection components
 $G_{n}\left(z\right)=-n!{\displaystyle\sum_{k=1,3,\ldots}\dfrac{z^{k}}{k^{n}\cdot k!}}$
@@
 each of which is itself an $E$-function, and these new functions arranged into a system of linear differential equations with coefficients in $\mathbb{Q}\left(z\right)$ and a unique finite singularity at $z=0$.
-By employing the Kolchin-Ostrowski differential-field framework (see Srinivasan [18]), we then can show that, for any $n\in\mathbb{Z}_{\geq 1}$, the $G_{0}\left(z\right),H_{0}\left(z\right),G_{1}\left(z\right),H_{1}\left(z\right),\ldots,G_{n}\left(z\right),H_{n}\left(z\right)$ are algebraically independent over $\mathbb{C}\left(z\right)$.
-This allows us to conclude, via the refined Siegel-Shidlovskii theorem (see Beukers [15]), that for $z=1$ the set of values $\left\{G_{0}\left(1\right),H_{0}\left(1\right),G_{1}\left(1\right),H_{1}\left(1\right),\ldots,G_{n}\left(1\right),H_{n}\left(1\right)\right\}$ are algebraically independent over $\overline{\mathbb{Q}}$, and therefore transcendental.
+By employing the Kolchin-Ostrowski differential-field framework (see Srinivasan [18]), we then can show that, for any $m\in\mathbb{Z}_{\geq 1}$, the functions $G_{1}\left(z\right),H_{1}\left(z\right),G_{2}\left(z\right),H_{2}\left(z\right),\ldots,G_{m}\left(z\right),H_{m}\left(z\right)$ are algebraically independent over $\mathbb{C}\left(z,G_{0}\left(z\right),H_{0}\left(z\right)\right)$.
+This allows us to conclude, via the refined Siegel-Shidlovskii theorem (see Beukers [15]), that for $z=1$ the set of values $\left\{G_{1}\left(1\right),H_{1}\left(1\right),G_{2}\left(1\right),H_{2}\left(1\right),\ldots,G_{m}\left(1\right),H_{m}\left(1\right)\right\}$ are algebraically independent over $\overline{\mathbb{Q}}\left(G_{0}\left(1\right),H_{0}\left(1\right)\right)$, and therefore transcendental.
 Finally, recognizing that
@@
 $n!\left(1-\dfrac{1}{2^{n+1}}+O\left(\dfrac{1}{3^{n}}\right)\right)-\left(-1\right)^{n+1}\left[W\left(n\right)\right]^{n}\exp\left(-\dfrac{n}{W\left(n\right)}\right)\sqrt{\dfrac{2\pi n}{W\left(n\right)+1}}\left(1+o\left(1\right)\right)$
+$n!\left(1-\dfrac{1}{2^{n+1}}+O\left(\dfrac{1}{3^{n}}\right)\right)+\left(n!\right)o\left(\dfrac{1}{3^{n}}\right)$
 $n!\left(1-\dfrac{1}{2^{n+1}}+O\left(\dfrac{1}{3^{n}}\right)\right).$
@@
 $\begin{array}[]{c}\widetilde{\delta}^{\left(n\right)}=e\left[-n!\left(1+\dfrac{1}{2^{n}\cdot 2!}+\dfrac{1}{3^{n}\cdot 3!}+\dfrac{1}{4^{n}\cdot 4!}+\cdots\right)-n!\left(1-\dfrac{1}{2^{n}\cdot 2!}+\dfrac{1}{3^{n}\cdot 3!}-\dfrac{1}{4^{n}\cdot 4!}+\cdots\right)\right.\\ \left.+\dfrac{1}{e}\left(-1\right)^{n+1}e{\displaystyle\int_{1}^{\infty}\left[\ln\left(u\right)\right]^{n}e^{-u}du}\right]\end{array}$
-$e\left[-2n!\left(1+\dfrac{1}{2\cdot 3^{n+1}}+O\left(\dfrac{1}{5^{n}}\right)\right)\right]$
+$e\left[-2n!\left(1+\dfrac{1}{2\cdot 3^{n+1}}+O\left(\dfrac{1}{5^{n}}\right)\right)+\left(n!\right)o\left(\dfrac{1}{5^{n}}\right)\right]$
 $n!\left(2e+\dfrac{e}{3^{n+1}}+O\left(\dfrac{1}{5^{n}}\right)\right).$

### v6

@@
-Arithmetic Properties of Several Generalized-Constant Sequences, with Implications for $\boldsymbol{\Gamma^{\left(n\right)}\left(1\right)}$
+Transcendence Results for $\boldsymbol{\Gamma^{\left(n\right)}\left(1\right)}$ and Related Generalized-Constant Sequences
 Abstract.
@@
 The two constants are related through a well-known equation of Hardy, equivalent to $\gamma+\delta/e=\textrm{Ein}\left(1\right)$, which recently has been generalized to $\gamma^{\left(n\right)}+\delta^{\left(n\right)}/e=\eta^{\left(n\right)},\>n\geq 0$ for sequences of constants $\gamma^{\left(n\right)}$, $\delta^{\left(n\right)}$, and $\eta^{\left(n\right)}$ (given respectively by raw, conditional, and partial moments of the Gumbel(0,1) probability distribution).
-Investigating the $\gamma^{\left(n\right)}$ through recurrence relations (where $\gamma^{\left(0\right)}=1$ and $\gamma^{\left(1\right)}=\gamma$), we find that at least one of the pair $\{\gamma,\gamma^{\left(2\right)}\}$ and at least two of each set $\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\}$, for $2\leq n\leq 26$, are transcendental, with analogous results for the sequence $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$.
+Investigating $\gamma^{\left(n\right)}=\left(-1\right)^{n}\Gamma^{\left(n\right)}\left(1\right),\>n\geq 1$ through recurrence relations, we find that the $\gamma^{\left(n\right)}$ are transcendental infinitely often and that at least two of each set $\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\},\>2\leq n\leq 26$ are transcendental (with analogous results for $\Gamma^{\left(n\right)}\left(1\right)$).
 We then show, via a theorem of Shidlovskii, that the $\eta^{\left(n\right)}$ are algebraically independent, and therefore transcendental, for all $n\geq 0$, implying that at least one of each pair, $\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e\}$ and $\{\gamma^{\left(n\right)},\delta^{\left(n\right)}\}$, and at least two of the triple $\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e,\delta^{\left(n\right)}\}$, are transcendental for all $n\geq 1$.
-Further analysis of the $\gamma^{\left(n\right)}$ and $\eta^{\left(n\right)}$ reveals that the values $\delta^{\left(n\right)}/e$ are transcendental infinitely often with positive asymptotic density.
+Further analysis of the $\gamma^{\left(n\right)}$ and $\eta^{\left(n\right)}$ reveals that both the $\delta^{\left(n\right)}/e$ and $\delta^{\left(n\right)}$ are transcendental infinitely often with asymptotic densities of at least $1/2$.
 Finally, we provide parallel results for the sequences $\widetilde{\delta}^{\left(n\right)}$ and $\widetilde{\eta}^{\left(n\right)}$ satisfying the “non-alternating analogue” equation $\gamma^{\left(n\right)}+\widetilde{\delta}^{\left(n\right)}/e=\widetilde{\eta}^{\left(n\right)}$.
 Key words and phrases:
-Euler-Mascheroni constant; Euler-Gompertz constant; Gumbel distribution; irrationality; transcendence; algebraic independence; gamma-function derivatives.
+Euler-Mascheroni constant; Euler-Gompertz constant; Gumbel distribution; transcendence; algebraic independence; gamma-function derivatives.
 2000 Mathematics Subject Classification:
@@
 so that (5) may be rewritten as
-$\gamma^{\left(n\right)}+\dfrac{\delta^{\left(n\right)}}{e}=\eta^{\left(n\right)}.$ (8)
+$\gamma^{\left(n\right)}+\dfrac{\delta^{\left(n\right)}}{e}=\eta^{\left(n\right)},$ (8)
+and note that
+$\gamma^{\left(n\right)}=\left(-1\right)^{n}\Gamma^{\left(n\right)}\left(1\right).$
 Further, we will refer to the sequences $\gamma^{\left(n\right)}$, $\delta^{\left(n\right)}$, and $\eta^{\left(n\right)}$ as the generalized Euler-Mascheroni, Euler-Gompertz, and Eta constants, respectively (with $\gamma^{\left(1\right)}=\gamma$, $\delta^{\left(1\right)}=\delta=-e\textrm{Ei}\left(-1\right)$, and $\eta^{\left(1\right)}=\eta=\textrm{Ein}\left(1\right)$ denoting the ordinary Euler-Mascheroni, Euler-Gompertz, and Eta constants).
@@
 99 $362,526.2891146549\ldots$ $6.3945118625\ldots$ $362,528.6415241055\ldots$
-$10$ $\qquad 3,627,042.4127568947\ldots\qquad$ $11.3803468877\ldots$ $\qquad 3,627,038.2261612415\ldots\qquad$
+$10$ $3,627,042.4127568947\ldots$ $11.3803468877\ldots$ $3,627,038.2261612415\ldots$
 $11$ $39,907,084.1514313358\ldots$ $20.9346984188\ldots$ $39,907,091.8528764918\ldots$
@@
 1.3. Outline of Results
-In Section 2, we begin our study of the arithmetic properties of the three generalized-constant sequences by investigating recurrence relations associated with the $\gamma^{\left(n\right)}$.
-These recurrences reveal that at least one element of the pair $\left\{\gamma,\gamma^{\left(2\right)}\right\}$ and at least two elements of each set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$, for $n\in\left\{2,3,\ldots,26\right\}$, are transcendental.
-As will be seen, the latter result follows from verification, for $n\in\left\{2,3,\ldots,26\right\}$, of a certain reasonable conjecture (which, if true for all $n\in\mathbb{Z}_{\geq 2}$, necessitates that the $\gamma^{\left(n\right)}$ are transcendental infinitely often).
-Since $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$, all conclusions for the $\gamma^{\left(n\right)}$ apply equally well to the sequence $\Gamma^{\left(n\right)}\left(1\right)$.
+In Section 2, we begin our study of the arithmetic properties of the three generalized-constant sequences by investigating recurrence relations associated with $\gamma^{\left(n\right)}=\left(-1\right)^{n}\Gamma^{\left(n\right)}\left(1\right)$.
+These recurrences reveal that the $\gamma^{\left(n\right)}$ are transcendental infinitely often and that at least two elements of each set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$, for $n\in\left\{2,3,\ldots,26\right\}$, are transcendental (with analogous implications for $\Gamma^{\left(n\right)}\left(1\right)$).
+As will be seen, the latter result follows from verification, for $n\in\left\{2,3,\ldots,26\right\}$, of a certain reasonable conjecture that, if true for all $n\in\mathbb{Z}_{\geq 2}$, enables one to derive an asymptotic lower bound on the density of transcendental terms.
 In Section 3, we employ a theorem of Shidlovskii (1989) to demonstrate the algebraic independence, and therefore transcendence, of the $\eta^{\left(n\right)}$ for all $n\geq 0$.
 This implies that at least one element of each pair, $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e\right\}$ and $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}\right\}$, and at least two elements of the triple $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e,\delta^{\left(n\right)}\right\}$, are transcendental for all $n\geq 1$.
-Moreover, further analysis of the $\gamma^{\left(n\right)}$ and $\eta^{\left(n\right)}$ shows that the values $\delta^{\left(n\right)}/e$ are transcendental infinitely often, with the density of the set of transcendental terms having asymptotic lower bound $1/2$.
+Moreover, further analysis of the $\gamma^{\left(n\right)}$ and $\eta^{\left(n\right)}$ shows that both the $\delta^{\left(n\right)}/e$ and $\delta^{\left(n\right)}$ are transcendental infinitely often, with the densities of both sets of transcendental terms having asymptotic lower bound $1/2$.
 Finally, Section 4 addresses two additional sequences, $\widetilde{\delta}^{\left(n\right)}$ and $\widetilde{\eta}^{\left(n\right)}$, satisfying the “non-alternating analogue” equation $\gamma^{\left(n\right)}+\widetilde{\delta}^{\left(n\right)}/e=\widetilde{\eta}^{\left(n\right)}$, and provides results parallel to those for the original system.
@@
 then follow from the complete exponential Bell polynomials $\textrm{E}_{X}\left[X^{n}\right]=B_{n}\left(\kappa_{1},\kappa_{2},\ldots,\kappa_{n}\right)$, with $\textrm{E}_{X}\left[X^{0}\right]=1$ and $\kappa_{1}=\textrm{E}_{X}\left[X\right]$.
-For the case of $X\equiv X_{\textrm{G}}\sim\textrm{Gumbel}\left(0,1\right)$, the identities in (9) and (10) can be simplified by setting
+For the case of $X\sim F_{X}\equiv\textrm{Gumbel}\left(0,1\right)$, the identities in (9) and (10) can be simplified by setting
 $\kappa_{\ell}=\begin{cases}\gamma,&\ell=1\\ \left(\ell-1\right)!\zeta\left(\ell\right),&\ell\in\left\{2,3,\ldots\right\}\end{cases},$
@@
 $\left(\ell-1\right)!\zeta\left(\ell\right)=\gamma^{\left(\ell\right)}-\gamma\cdot\gamma^{\left(\ell-1\right)}-{\displaystyle\sum_{i=1}^{\ell-2}}\dfrac{\left(\ell-1\right)!}{\left(\ell-i-1\right)!}\zeta\left(i+1\right)\gamma^{\left(\ell-i-1\right)},\quad\ell\in\mathbb{Z}_{\geq 2}.$ (12)
-Further, we can rewrite (11) as:
+Further, we can iterate (11) to find:
 $\begin{array}[]{c}\gamma^{\left(1\right)}=\gamma,\\ \gamma^{\left(2\right)}=\gamma^{2}+\zeta\left(2\right),\\ \gamma^{\left(3\right)}=\gamma^{3}+3\zeta\left(2\right)\gamma+2\zeta\left(3\right),\\ \gamma^{\left(4\right)}=\gamma^{4}+6\zeta\left(2\right)\gamma^{2}+8\zeta\left(3\right)\gamma+\dfrac{27}{2}\zeta\left(4\right),\\ \cdots;\end{array}$ (13)
@@
 Lemma 2.
-The variety $\mathcal{V}_{n}$ is irreducible and has function field of transcendence degree $n+1$.
+The variety $\mathcal{V}_{n}$ is irreducible and has a function field of transcendence degree $n+1$.
 ?proofname?.
@@
 According to the Jacobian criterion for algebraic independence (which is applicable because we are working over a field of characteristic 0), the $n+1$ polynomials corresponding to $\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}$, for any $\theta\in\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)$, are algebraically independent if and only if the determinant of the $\left(n+1\right)\times\left(n+1\right)$ Jacobian matrix
-$\mathbf{J}_{-\theta}^{\left(n\right)}\left(j,\ell\right)=\left[\frac{d\gamma^{\left(j\right)}}{d\kappa_{\ell}}\right],\quad j\in\left(\left\{1,n,n+1,\ldots,2n\right\}\setminus\left\{k:\theta=\gamma^{\left(k\right)}\right\}\right),\ell\in\left\{1,2,3,5,\ldots,2n-1\right\}$
+$\mathbf{J}_{-\theta}^{\left(n\right)}=\left[\frac{d\gamma^{\left(j\right)}}{d\kappa_{\ell}}\right],\quad j\in\left(\left\{1,n,n+1,\ldots,2n\right\}\setminus\left\{k:\theta=\gamma^{\left(k\right)}\right\}\right),\ell\in\left\{1,2,3,5,\ldots,2n-1\right\}$
 is not identically zero. To determine whether or not this condition holds, first note that
@@
 Specializing the Jacobian at the convenient realization $\mathcal{K}_{n}^{*}=\left\{\kappa_{1}=1,\kappa_{2}=0,\kappa_{3}=0,\kappa_{5}=0,\ldots,\kappa_{2n-1}=0\right\}$, for which $mb_{m}\kappa_{2}^{m-1}=0$, then gives
-$\left.\mathbf{J}_{-\theta}^{\left(n\right)}\left(j,\ell\right)\right|_{\mathcal{K}_{n}^{*}}=\left.\left[\frac{\partial\gamma^{\left(j\right)}}{\partial\kappa_{\ell}}\right]\right|_{\mathcal{K}_{n}^{*}},$
+$\left.\mathbf{J}_{-\theta}^{\left(n\right)}\right|_{\mathcal{K}_{n}^{*}}=\left.\left[\frac{\partial\gamma^{\left(j\right)}}{\partial\kappa_{\ell}}\right]\right|_{\mathcal{K}_{n}^{*}},$
 which simplifies further to
-$\left.\mathbf{J}_{-\theta}^{\left(n\right)}\left(j,\ell\right)\right|_{\mathcal{K}_{n}^{*}}=\binom{j}{\ell}$
+$\left.\mathbf{J}_{-\theta}^{\left(n\right)}\right|_{\mathcal{K}_{n}^{*}}=\left[\binom{j}{\ell}\right$]
 because $\partial\gamma^{\left(j\right)}/\partial\kappa_{\ell}=\binom{j}{\ell}\gamma^{\left(j-\ell\right)}$ (from the Bell-polynomial identity $\partial B_{j}\left(\kappa_{1},\kappa_{2},\ldots,\kappa_{j}\right)=$ $\binom{j}{\ell}B_{j-\ell}\left(\kappa_{1},\kappa_{2},\ldots,\kappa_{j-\ell}\right)$) and $\gamma^{\left(j-\ell\right)}=\kappa_{1}^{j-\ell}=1$ at $\mathcal{K}_{n}^{*}$.
@@
 where $\mathbf{M}_{-\theta}^{\left(n\right)}$ is the $\left(n+1\right)\times\left(n+1\right)$ matrix
-$\mathbf{M}_{-\theta}^{\left(n\right)}\left(j,\ell\right)=\binom{j}{\ell},\quad j\in\left(\left\{1,n,n+1,\ldots,2n\right\}\setminus\left\{k:\theta=\gamma^{\left(k\right)}\right\}\right),\ell\in\left\{1,2,3,5,\ldots,2n-1\right\}.$
-To see that $\det\left(\mathbf{M}_{-\theta}^{\left(n\right)}\right)\neq 0$, we note that $\mathbf{M}_{-\theta}^{\left(n\right)}$ is a submatrix of the lower-triangular Pascal matrix,
-$\mathbf{P}^{\left(n\right)}\left(r,c\right)=\binom{r}{c},\quad r\in\left\{0,1,\ldots,2n\right\},c\in\left\{0,1,\ldots,2n\right\},$
+$\mathbf{M}_{-\theta}^{\left(n\right)}=\left[\binom{j}{\ell}\right],\quad j\in\left(\left\{1,n,n+1,\ldots,2n\right\}\setminus\left\{k:\theta=\gamma^{\left(k\right)}\right\}\right),\ell\in\left\{1,2,3,5,\ldots,2n-1\right\}.$
+To see that $\det\left(\mathbf{M}_{-\theta}^{\left(n\right)}\right)\neq 0$, we note that $\mathbf{M}_{-\theta}^{\left(n\right)}$ is a submatrix of the $\left(2n+1\right)\times\left(2n+1\right)$ lower-triangular Pascal matrix,
+$\mathbf{P}^{\left(n\right)}=\left[\binom{r}{c}\right],\quad r\in\left\{0,1,\ldots,2n\right\},c\in\left\{0,1,\ldots,2n\right\},$
 which satisfies the conditions for a lower strictly totally positive (LSTP) matrix given by Section 2.3, Theorem 2.8 of Pinkus [8].
@@
 2.3. Transcendence Results
-The transformation of the original polynomials of (16) into the reduced form of (17) lays the foundation for part (ii) of Theorem 1 below.
+The transformation of the original polynomials of (16) into the reduced form of (17) lays the foundation for parts (iii) and (iv) of Theorem 1 below.
 However, one additional condition on the $\widetilde{P}_{n,\theta}$ is required.
@@
 Now define the $\left(m-1\right)\times\left(m-1\right)$ Jacobian matrix
-$\mathbf{J}_{m,\theta}^{\left(n\right)}\left(i,j\right)=\left[\frac{\partial P_{i}}{\partial\gamma^{\left(j\right)}}\right],\quad i\in\left\{2,3,\ldots,m\right\},j\in\left(\left\{2,3,\ldots,m-1\right\}\cup\left\{k:\theta=\gamma^{\left(k\right)}\right\}\right),$
+$\mathbf{J}_{m,\theta}^{\left(n\right)}=\left[\frac{\partial P_{i}}{\partial\gamma^{\left(j\right)}}\right],\quad i\in\left\{2,3,\ldots,m\right\},j\in\left(\left\{2,3,\ldots,m-1\right\}\cup\left\{k:\theta=\gamma^{\left(k\right)}\right\}\right),$
 for which $\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)\in\mathcal{R}_{n}.$
@@
 For fixed $n\in\mathbb{Z}_{\geq 2}$, $m\in\left\{2,3,\ldots,n\right\}$, and $\theta\in\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)$, if $\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)\notin\mathcal{J}_{n,\boldsymbol{\gamma^{*}}}$ – or equivalently, if
-$\left.\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)\right|_{\boldsymbol{\gamma^{*}}}\neq 0$ (21)
+$\det\left(\left.\mathbf{J}_{m,\theta}^{\left(n\right)}\right|_{\gamma^{*}}\right)\neq 0$ (21)
 – then there exists a polynomial $\widetilde{P}_{m,\theta}\in\mathbb{Z}\left[\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right]$ satisfying (17) such that, when written in the form of (19), at least one coefficient $A_{d}^{\left(m,\theta\right)}\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right),\>d\geq 1$ must remain non-zero at $\boldsymbol{\gamma^{*}}$.
@@
 Rewriting each $b_{d}^{\left(m,\theta\right)}\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right)$ as $\beta_{d}^{\left(m,\theta\right)}\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right)/u^{q_{d}^{\left(m,\theta\right)}}$ with $\beta_{d}^{\left(m,\theta\right)}\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right)\in\mathcal{U}_{m,\theta}$ and $q_{d}^{\left(m,\theta\right)}\in\mathbb{Z}_{\geq 0}$, we then can multiply $M_{m,\theta}\left(t\right)$ by $u^{\widehat{q}^{\left(m,\theta\right)}}$, where $\widehat{q}^{\left(m,\theta\right)}=\underset{d}{\max}\left\{q_{d}^{\left(m,\theta\right)}\right\}$, to obtain
-$\begin{array}[]{c}M_{m,\theta}^{*}\left(t\right)=u^{\widehat{q}^{\left(m,\theta\right)}}t^{D}+\beta_{D-1}^{\left(m,\theta\right)}u^{\left(\widehat{q}^{\left(m,\theta\right)}-q_{D-1}^{\left(m,\theta\right)}\right)}\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right)t^{D-1}+\cdots\\ +\beta_{0}^{\left(m,\theta\right)}u^{\left(\widehat{q}^{\left(m,\theta\right)}-q_{0}^{\left(m,\theta\right)}\right)}\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right)\in\mathcal{U}_{m,\theta}\left[t\right],\end{array}$
-with $M_{m,\theta}^{*}\left(\theta\right)=0$ in $\left(\mathcal{S}_{m}\right)_{s}$ and leading coefficient $u^{\widehat{q}^{\left(m,\theta\right)}}$.
-Since this equality holds in the localization $\left(\mathcal{S}_{m}\right)_{s}$, there exists $q\in\mathbb{Z}_{\geq 0}$ such that $s^{q}M_{m,\theta}^{*}\left(\theta\right)=0$ in $\mathcal{S}_{m}$.
+$\begin{array}[]{c}\widehat{M}_{m,\theta}\left(t\right)=u^{\widehat{q}^{\left(m,\theta\right)}}t^{D}+\beta_{D-1}^{\left(m,\theta\right)}u^{\left(\widehat{q}^{\left(m,\theta\right)}-q_{D-1}^{\left(m,\theta\right)}\right)}\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right)t^{D-1}+\cdots\\ +\beta_{0}^{\left(m,\theta\right)}u^{\left(\widehat{q}^{\left(m,\theta\right)}-q_{0}^{\left(m,\theta\right)}\right)}\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right)\in\mathcal{U}_{m,\theta}\left[t\right],\end{array}$
+with $\widehat{M}_{m,\theta}\left(\theta\right)=0$ in $\left(\mathcal{S}_{m}\right)_{s}$ and leading coefficient $u^{\widehat{q}^{\left(m,\theta\right)}}$.
+Since this equality holds in the localization $\left(\mathcal{S}_{m}\right)_{s}$, there exists $q\in\mathbb{Z}_{\geq 0}$ such that $s^{q}\widehat{M}_{m,\theta}\left(\theta\right)=0$ in $\mathcal{S}_{m}$.
 Now specialize at $\boldsymbol{\gamma^{*}}$, and note that $u\notin\mathcal{T}_{m,\theta}$ and $s\notin\mathcal{\overline{J}}_{m,\boldsymbol{\gamma^{*}}}$ together force $\left.u\right|_{\boldsymbol{\gamma^{*}}}\neq 0$ and $\left.s\right|_{\boldsymbol{\gamma^{*}}}\neq 0$.
@@
 First, we see that
-$\left.\left(s^{q}M_{m,\theta}^{*}\left(\theta\right)\right)\right|_{\boldsymbol{\gamma^{*}}}=0\Longrightarrow\left(\left.s\right|_{\boldsymbol{\gamma^{*}}}\right)^{q}M_{m,\theta}^{*}\left(\left.\theta\right|_{\boldsymbol{\gamma^{*}}}\right)=0$
-$\Longrightarrow M_{m,\theta}^{*}\left(\left.\theta\right|_{\boldsymbol{\gamma^{*}}}\right)=0;$
-and second, for the leading coefficient of $M_{m,\theta}^{*}\left(t\right)$, we find
+$\left.\left(s^{q}\widehat{M}_{m,\theta}\left(\theta\right)\right)\right|_{\boldsymbol{\gamma^{*}}}=0\Longrightarrow\left(\left.s\right|_{\boldsymbol{\gamma^{*}}}\right)^{q}\widehat{M}_{m,\theta}\left(\left.\theta\right|_{\boldsymbol{\gamma^{*}}}\right)=0$
+$\Longrightarrow\widehat{M}_{m,\theta}\left(\left.\theta\right|_{\boldsymbol{\gamma^{*}}}\right)=0;$
+and second, for the leading coefficient of $\widehat{M}_{m,\theta}\left(t\right)$, we find
 $\left.u^{\widehat{q}^{\left(m,\theta\right)}}\right|_{\boldsymbol{\gamma^{*}}}=\left(\left.u\right|_{\boldsymbol{\gamma^{*}}}\right)^{\widehat{q}^{\left(m,\theta\right)}}\neq 0.$
-These facts imply that $M_{m,\theta}^{*}\left(t\right)$ specializes to a non-zero polynomial in $t$ that vanishes at $t=\left.\theta\right|_{\boldsymbol{\gamma^{*}}}$, and so cannot be a constant.
+These facts imply that $\widehat{M}_{m,\theta}\left(t\right)$ specializes to a non-zero polynomial in $t$ that vanishes at $t=\left.\theta\right|_{\boldsymbol{\gamma^{*}}}$, and so cannot be a constant.
 Thus, at least one coefficient of the polynomial, for degree $d\geq 1$, must remain non-zero at $\boldsymbol{\gamma^{*}}$.
-Finally, let $\widehat{M}_{m,\theta}^{*}\left(t\right)$ be a primitive integer multiple of $M_{m,\theta}^{*}\left(t\right)$ (viewed in $\mathbb{Z}\left[\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right]\left[t\right]$), and factor $\widehat{M}_{m,\theta}^{*}\left(t\right)$ into irreducibles in this unique factorization domain.
-We then can choose an irreducible factor $\widetilde{P}_{m,\theta}\left(t\right)$ such that $\widetilde{P}_{m,\theta}\left(\left.\theta\right|_{\boldsymbol{\gamma^{*}}}\right)=0$ and $\left.\widetilde{P}_{m,\theta}\left(t\right)\right|_{\boldsymbol{\gamma^{*}}}\not\equiv 0$ (because $\left.\widehat{M}_{m,\theta}^{*}\left(t\right)\right|_{\boldsymbol{\gamma^{*}}}$ is a non-zero polynomial with root $\left.\theta\right|_{\boldsymbol{\gamma^{*}}}$).
+Finally, let $\widehat{M}_{m,\theta}^{\prime}\left(t\right)$ be a primitive integer multiple of $\widehat{M}_{m,\theta}\left(t\right)$ (viewed in $\mathbb{Z}\left[\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right]\left[t\right]$), and factor $\widehat{M}_{m,\theta}^{\prime}\left(t\right)$ into irreducibles in this unique factorization domain.
+We then can choose an irreducible factor $\widetilde{P}_{m,\theta}\left(t\right)$ such that $\widetilde{P}_{m,\theta}\left(\left.\theta\right|_{\boldsymbol{\gamma^{*}}}\right)=0$ and $\left.\widetilde{P}_{m,\theta}\left(t\right)\right|_{\boldsymbol{\gamma^{*}}}\not\equiv 0$ (because $\left.\widehat{M}_{m,\theta}^{\prime}\left(t\right)\right|_{\boldsymbol{\gamma^{*}}}$ is a non-zero polynomial with root $\left.\theta\right|_{\boldsymbol{\gamma^{*}}}$).
 This $\widetilde{P}_{m,\theta}\in\mathbb{Z}\left[\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right]$ not only provides the non-trivial polynomial relation in $\theta$ required by (17), but also has the property that when written as a polynomial in $\theta$, it possesses at least one coefficient, for degree $d\geq 1$, that remains non-zero at $\boldsymbol{\gamma^{*}}$.
@@
 [^9] For each $n$, we executed a Python program (on an Apple M4 Max MacBook computer with 48 GB memory) to provide rigorous certification of the sufficient condition in (21).
-The program first evaluated $\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)$ symbolically for each $m\in\left\{2,3,\ldots,n\right\}$, and then confirmed $\left.\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)\right|_{\boldsymbol{\gamma^{*}}}\neq 0$ using interval arithmetic.
+The program first evaluated $\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)$ symbolically for each $m\in\left\{2,3,\ldots,n\right\}$, and then confirmed $\det\left(\left.\mathbf{J}_{m,\theta}^{\left(n\right)}\right|_{\gamma^{*}}\right)\neq 0$ using interval arithmetic.
 The maximum value of $n$ tested (i.e., $26$) was simply the largest $n$ for which the program completed all checks within approximately 24 hours.
@@
 For fixed $n\in\mathbb{Z}_{\geq 2}$,
-$\left.\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)\right|_{\boldsymbol{\gamma^{*}}}\neq 0$
+$\det\left(\left.\mathbf{J}_{m,\theta}^{\left(n\right)}\right|_{\gamma^{*}}\right)\neq 0$
 for all $m\in\left\{2,3,\ldots,n\right\}$ and $\theta\in\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)$.
@@
 (i) at least one element of the pair $\left\{\gamma,\gamma^{\left(2\right)}\right\}$ is transcendental;
-(ii) if Conjecture 1 holds for a given $n\in\mathbb{Z}_{\geq 2}$, then at least two elements of the set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$ are transcendental;
-(iii) if Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, then $\gamma^{\left(n\right)}$ is transcendental infinitely often; and
+(ii) $\gamma^{\left(n\right)}$ is transcendental infinitely often;
+(iii) if Conjecture 1 holds for a given $n\in\mathbb{Z}_{\geq 2}$, then at least two elements of the set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$ are transcendental; and
 (iv) if Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, then the density of the set of transcendental $\gamma^{\left(n\right)}$ evaluated at $N\in\mathbb{Z}_{\geq 2}$ (i.e., $\left\{n\leq N:\gamma^{\left(n\right)}\textrm{ is transcendental}\right\}/N$) has asymptotic lower bound $\Omega\left(\ln\left(N\right)/N\right)$ as $N\rightarrow\infty$.
@@
 (i) Since $\zeta\left(2\right)=\pi^{2}/6$ is transcendental, one can see from the first two lines of (13) that if either $\gamma$ or $\gamma^{\left(2\right)}=\gamma^{2}+\zeta\left(2\right)$ is algebraic, then the other member of the pair must be transcendental.
-(ii) We proceed by induction on $m\in\left\{2,3,\ldots,n\right\}$.
+(ii) For $X\sim\textrm{Gumbel}\left(0,1\right)$, the moment-generating function, $\textrm{M}_{X}\left(t\right)=\textrm{E}_{X}\left[e^{tX}\right],\>t<1$, is given by
+$\Gamma\left(1-t\right)={\displaystyle\sum_{n=0}^{\infty}}\textrm{E}_{X}\left[X^{n}\right]\dfrac{t^{n}}{n!}={\displaystyle\sum_{n=0}^{\infty}}\gamma^{\left(n\right)}\dfrac{t^{n}}{n!}.$
+Euler’s reflection formula then yields
+$\Gamma\left(1+t\right)\Gamma\left(1-t\right)=\dfrac{\pi t}{\sin\left(\pi t\right)}$
+$\Longrightarrow\left({\displaystyle\sum_{m=0}^{\infty}}\gamma^{\left(m\right)}\dfrac{\left(-t\right)^{m}}{m!}\right)\left({\displaystyle\sum_{n=0}^{\infty}}\gamma^{\left(n\right)}\dfrac{t^{n}}{n!}\right)={\displaystyle\sum_{k=0}^{\infty}}c_{k}\pi^{2k}t^{2k},$
+where $c_{k}\in\mathbb{Q}^{\times}$ for all $k\geq 0$, and we can compare the coefficient of $t^{2k}$ on both sides of the above equation to obtain
+${\displaystyle\sum_{j=0}^{2k}}\dfrac{\left(-1\right)^{j}\gamma^{\left(j\right)}\gamma^{\left(2k-j\right)}}{j!\left(2k-j\right)!}=c_{k}\pi^{2k}$ (22)
+for all $k$.
+Suppose (for purposes of contradiction) that only finitely many $\gamma^{\left(n\right)}$ are transcendental, and choose $n^{*}\in\mathbb{Z}_{\geq 0}$ such that $\gamma^{\left(n\right)}\in\overline{\mathbb{Q}}$ for all $n>n^{*}$, with $\mathcal{W}=\textrm{span}_{\overline{\mathbb{Q}}}\left\{\gamma^{\left(0\right)},\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\ldots,\gamma^{\left(n^{*}\right)}\right\}$ denoting a finite-dimensional $\overline{\mathbb{Q}}$-vector space.
+Now note that for any $k>n^{*}$, each summand on the left-hand side of (22) has at least one algebraic factor $\gamma^{\left(i_{k}\right)}$, for some $i_{k}>n^{*}$ (because $j+\left(2k-j\right)=2k>2n^{*}\Longrightarrow\left(j>n^{*}\right)\vee\left(2k-j>n^{*}\right)$).
+It then follows that every summand lies in $\mathcal{W}$, and therefore that $c_{k}\pi^{2k}\in\mathcal{W}$ for all $k>n^{*}$.
+Since $c_{k}\in\mathbb{Q}^{\times}$, we see that $\pi^{2k}\in\mathcal{W}$ for all $k>n^{*}$, and thus that all $n^{*}+2$ elements $\left\{\pi^{2\left(n^{*}+1\right)},\pi^{2\left(n^{*}+2\right)},\ldots,\pi^{2\left(2n^{*}+2\right)}\right\}$ must lie in the space $\mathcal{W}$.
+However, the dimension of $\mathcal{W}$ is at most $n^{*}+1$, forcing a non-trivial $\overline{\mathbb{Q}}$-linear dependence among these $n^{*}+2$ elements; that is, a non-trivial polynomial relation over $\overline{\mathbb{Q}}$ satisfied by $\pi^{2}$, which contradicts the transcendence of $\pi$.
+(iii) We proceed by induction on $m\in\left\{2,3,\ldots,n\right\}$.
 For the base case, $m=2$, it is known from part (i) that $\mathcal{G}_{2}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}=\left\{\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\gamma^{\left(4\right)}\right\}$ contains at least one transcendental element.
@@
 (It is important to note that this approach cannot be extended to show that “at least three” elements of a set of distinct generalized Euler-Mascheroni constants are transcendental because the presence of a second transcendental constant removes the necessity that, for a given $\theta$, all coefficients of $\widetilde{P}_{m,\theta}\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)=0$ are algebraic.)
-(iii) Assuming Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, part (ii) implies that the set $\mathcal{G}_{n}^{\left(\textrm{U}\right)}$ contains at least one transcendental element for all $n$.
-Since $n$ can be arbitrarily large, this precludes the possibility that there exists a finite $n^{*}\in\mathbb{Z}_{\geq 1}$ such that $\gamma^{\left(n\right)}$ is algebraic for all $n\geq n^{*}$.
-(iv) As previously observed, if Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, then part (ii) forces each set $\mathcal{G}_{n}^{\left(\textrm{U}\right)}$ to contain at least one transcendental element $\gamma^{\left(k_{n}\right)},\>k_{n}\in\left\{n,n+1,\ldots,2n\right\}$.
-Taking $m=3^{i}$ for $i\in\mathbb{Z}_{\geq 1}$, we now consider the family of disjoint intervals $\left[3^{i},2\cdot 3^{i}\right]\ni k_{m}$ and note that we can choose $N\in\mathbb{Z}_{\geq 2}$ such that $N\geq 2\cdot 3^{i}=2m$ and the $k_{m}\in\left[3^{i},2\cdot 3^{i}\right]\subseteq\left\{1,2,\ldots,N\right\}$ are distinct.
+(iv) Assuming Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, part (iii) implies that the set $\mathcal{G}_{n}^{\left(\textrm{U}\right)}$ contains at least one transcendental element $\gamma^{\left(k_{n}\right)},\>k_{n}\in\left\{n,n+1,\ldots,2n\right\}$ for all $n$.
+Taking $m=3^{\ell}$ for $\ell\in\mathbb{Z}_{\geq 1}$, we now consider the family of disjoint intervals $\left[3^{\ell},2\cdot 3^{\ell}\right]\ni k_{m}$ and note that we can choose $N\in\mathbb{Z}_{\geq 2}$ such that $N\geq 2\cdot 3^{\ell}=2m$ and the $k_{m}\in\left[3^{\ell},2\cdot 3^{\ell}\right]\subseteq\left\{1,2,\ldots,N\right\}$ are distinct.
 This implies that the density of the set of transcendental $\gamma^{\left(n\right)}$ evaluated at $N$ is given by
@@
 with asymptotic lower bound $\Omega\left(\ln\left(N\right)/N\right)$ as $N\rightarrow\infty$. ∎
-Since Conjecture 1 has been verified computationally for $n\in\left\{2,3,\ldots,26\right\}$, we know from part (ii) of the above theorem that at least two elements of each associated set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$ are transcendental.
-Furthermore, it is easy to see from the $\textrm{Gumbel}\left(0,1\right)$ moment-generating function, $\textrm{M}_{X}\left(t\right)=\Gamma\left(1-t\right),\>t<1$, that
-$\gamma^{\left(n\right)}=\textrm{E}_{X}\left[X^{n}\right]=\left(-1\right)^{n}\Gamma^{\left(n\right)}\left(1\right),$
-implying that all four parts of Theorem 1 are equally valid if one replaces $\gamma^{\left(\bullet\right)}$ with $\Gamma^{\left(\bullet\right)}\left(1\right)$ everywhere.
-In particular: (a) at least two elements of each set $\left\{\Gamma^{\left(1\right)}\left(1\right),\Gamma^{\left(n\right)}\left(1\right),\Gamma^{\left(n+1\right)}\left(1\right),\ldots,\Gamma^{\left(2n\right)}\left(1\right)\right\}$ are transcendental for $n\in\left\{2,3,\ldots,26\right\}$; and (b) if Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, then $\Gamma^{\left(n\right)}\left(1\right)$ must be transcendental infinitely often, a result consistent with the common expectation that $\Gamma^{\left(n\right)}\left(1\right)$ is transcendental for all $n\geq 1$ (see, e.g., Rivoal [10] and Fischler and Rivoal [11]).
+Since Conjecture 1 has been verified computationally for $n\in\left\{2,3,\ldots,26\right\}$, we know from part (iii) of the above theorem that at least two elements of each associated set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$ are transcendental.
+Furthermore, the identity $\gamma^{\left(n\right)}=\left(-1\right)^{n}\Gamma^{\left(n\right)}\left(1\right)$ implies that all four parts of Theorem 1 are equally valid if one replaces $\gamma^{\left(\bullet\right)}$ with $\Gamma^{\left(\bullet\right)}\left(1\right)$ everywhere.
+In particular: (a) at least two elements of each set $\left\{\Gamma^{\left(1\right)}\left(1\right),\Gamma^{\left(n\right)}\left(1\right),\Gamma^{\left(n+1\right)}\left(1\right),\ldots,\Gamma^{\left(2n\right)}\left(1\right)\right\}$ are transcendental for $n\in\left\{2,3,\ldots,26\right\}$; and (b) $\Gamma^{\left(n\right)}\left(1\right)$ is transcendental infinitely often, a result consistent with the common expectation that $\Gamma^{\left(n\right)}\left(1\right)$ is transcendental for all $n\geq 1$ (see, e.g., Rivoal [10] and Fischler and Rivoal [11]).
 [^10] It may be of interest to some readers to note that the asymptotic lower bound on the density of the set of posited transcendental $\Gamma^{\left(n\right)}\left(1\right)$ is smaller (in magnitude) than the asymptotic lower bound on the density of the set of irrational odd-$n$ $\zeta\left(n\right)$ given by Fischler, Sprang, and Zudilin [12]: $\Omega\left(\dfrac{2^{\left(1-\varepsilon\right)\ln\left(N\right)/\ln\left(\ln\left(N\right)\right)}}{N/2}\right)=\Omega\left(N^{\left(1-\varepsilon\right)\ln\left(2\right)/\ln\left(\ln\left(N\right)\right)-1}\right),$ for arbitrarily small $\varepsilon>0$.
@@
 This allows us to make the following assertions:
-(I) The $F_{n}\left(z\right)$, like the $w_{n}\left(z\right)$, are $E$-functions (in the sense of Siegel) that satisfy an $\left(n+1\right)$-order non-homogeneous linear differential equation with coefficients in $\mathbb{Q}\left(z\right)$ and unique finite singularity at $z=0$ (see Chapter 7, Section 1 of Shidlovskii [14]).
+(I) The $F_{n}\left(z\right)$, like the $w_{n}\left(z\right)$, are $E$-functions (in the sense of Siegel) that satisfy an $\left(n+1\right)$-order non-homogeneous linear differential equation with coefficients in $\mathbb{Q}\left(z\right)$ and a unique finite singularity at $z=0$ (see Chapter 7, Section 1 of Shidlovskii [14]).
 (II) For any $m\in\mathbb{Z}_{\geq 1}$ and $z\in\overline{\mathbb{Q}}^{\times}$, the set of values $\left\{F_{0}\left(z\right),F_{1}\left(z\right),\ldots,F_{m}\left(z\right)\right\}$, like the set of values $\left\{w_{0}\left(z\right),w_{1}\left(z\right),\ldots,w_{m}\left(z\right)\right\}$, are algebraically independent over $\overline{\mathbb{Q}}$ (see Chapter 7, Section 1, Theorem 1 of Shidlovskii [14]; the theorem is restated and proved in Chapter 8, Section 3 of the same volume).
@@
 Theorem 3.
-For the sequence $\left\{\delta^{\left(n\right)}/e\right\}_{n\geq 1}$:
-(i) $\delta^{\left(n\right)}/e$ is transcendental infinitely often; and
-(ii) the density of the set of transcendental $\delta^{\left(n\right)}/e$ (i.e., $\left\{n\leq N:\delta^{\left(n\right)}/e\textrm{ is transcendental}\right\}/N$) has asymptotic lower bound $1/2$ as $N\rightarrow\infty$.
+For the sequences $\left\{\delta^{\left(n\right)}/e\right\}_{n\geq 1}$ and $\left\{\delta^{\left(n\right)}\right\}_{n\geq 1}$:
+(i) $\delta^{\left(n\right)}/e$ is transcendental infinitely often;
+(ii) the density of the set of transcendental $\delta^{\left(n\right)}/e$ (i.e., $\left\{n\leq N:\delta^{\left(n\right)}/e\textrm{ is transcendental}\right\}/N$) has asymptotic lower bound $1/2$ as $N\rightarrow\infty$;
+(iii) $\delta^{\left(n\right)}$ is transcendental infinitely often; and
+(iv) the density of the set of transcendental $\delta^{\left(n\right)}$ (i.e., $\left\{n\leq N:\delta^{\left(n\right)}\textrm{ is transcendental}\right\}/N$) has asymptotic lower bound $1/2$ as $N\rightarrow\infty$.
 ?proofname?.
@@
 $\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{G}\right)}\right)\leq n-\lfloor n/2\rfloor+1$ (24)
-for all $n$, and begin by noting that this inequality holds for $n<4$ because $\mathcal{F}_{n}^{\left(\mathcal{G}\right)}$ is generated by at most $|\mathcal{G}_{n}|$ elements, and $|\mathcal{G}_{n}|\leq n\leq n-\lfloor n/2\rfloor+1$ for $n\in\left\{1,2,3\right\}$.
+for all $n$, and begin by noting that this inequality holds for $n<4$ because $\mathcal{F}_{n}^{\left(\mathcal{G}\right)}$ is generated by at most $|\mathcal{G}_{n}|$ elements, and $|\mathcal{G}_{n}|=n\leq n-\lfloor n/2\rfloor+1$ for $n\in\left\{1,2,3\right\}$.
 In the following analysis, we treat the case of $n\geq 4$.
@@
 Given that $\mathcal{B}_{n}$ contains $n-\lfloor n/2\rfloor+1$ elements, the transcendence degree of $\mathcal{F}_{n}^{\left(\mathcal{G}\right)}$ (which is generated by $\mathcal{B}_{n}$) satisfies (24).
-Then, from the relation in (8), one can see that $\mathcal{F}_{n}^{\left(\mathcal{H}\right)}$ is contained in the composite field $\mathcal{F}_{n}^{\left(\mathcal{G}\right)}\left(\mathcal{F}_{n}^{\left(\mathcal{D}\right)}\right)$, implying
+Then, from the relation in (8), one can see that
+$\mathcal{F}_{n}^{\left(\mathcal{H}\right)}\subseteq\mathcal{F}_{n}^{\left(\mathcal{G}\right)}\left(\mathcal{F}_{n}^{\left(\mathcal{D}\right)}\right),$
+implying
 $\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{H}\right)}\right)\leq\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{G}\right)}\left(\mathcal{F}_{n}^{\left(\mathcal{D}\right)}\right)\right)\leq\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{G}\right)}\right)+\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{D}\right)}\right).$ (25)
@@
 $\underset{N\rightarrow\infty}{\liminf}\left(\dfrac{\lfloor N/2\rfloor-1}{N}\right)=\dfrac{1}{2}.$
+(iii) Retain $\mathcal{G}_{n}$ and $\mathcal{F}_{n}^{\left(\mathcal{G}\right)}$ as defined in the proof of part (i), and recall from Theorem 2 that $\eta^{\left(n\right)}=\gamma^{\left(n\right)}+\delta^{\left(n\right)}/e$ is algebraically independent of $\eta^{\left(0\right)}=1-1/e$ for all $n\in\mathbb{Z}_{\geq 1}$.
+Writing
+$\eta^{\left(n\right)}=\gamma^{\left(n\right)}+\delta^{\left(n\right)}\left(1-\eta^{\left(0\right)}\right)$ (26)
+as in the proof of Corollary 1(ii), further define $\mathcal{H}_{n}^{\prime}=\left\{\eta^{\left(0\right)},\eta,\eta^{\left(2\right)},\eta^{\left(3\right)},\ldots,\eta^{\left(n\right)}\right\}$ and $\mathcal{D}_{n}^{\prime}=\left\{\delta,\delta^{\left(2\right)},\delta^{\left(3\right)},\ldots,\delta^{\left(n\right)}\right\}$ for $n\in\mathbb{Z}_{\geq 1}$, where $\mathcal{F}_{n}^{\left(\mathcal{H}^{\prime}\right)}=\mathbb{Q}\left(\mathcal{H}_{n}^{\prime}\right)$ and $\mathcal{F}_{n}^{\left(\mathcal{D}^{\prime}\right)}=\mathbb{Q}\left(\mathcal{D}_{n}^{\prime}\right)$ denote the corresponding fields over $\mathbb{Q}$.
+From (26), we know that
+$\mathcal{F}_{n}^{\left(\mathcal{H}^{\prime}\right)}\subseteq\mathbb{Q}\left(\eta^{\left(0\right)},\mathcal{G}_{n}\right)\left(\mathcal{F}_{n}^{\left(\mathcal{D}^{\prime}\right)}\right),$
+which implies
+$\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{H}^{\prime}\right)}\right)\leq\mathrm{trdeg}\left(\mathbb{Q}\left(\eta^{\left(0\right)},\mathcal{G}_{n}\right)\left(\mathcal{F}_{n}^{\left(\mathcal{D}^{\prime}\right)}\right)\right)\leq\mathrm{trdeg}\left(\mathbb{Q}\left(\eta^{\left(0\right)},\mathcal{G}_{n}\right)\right)+\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{D}^{\prime}\right)}\right).$
+Then, since adjoining a single element to a field can increase the transcendence degree by at most one, we have
+$\mathrm{trdeg}\left(\mathbb{Q}\left(\eta^{\left(0\right)},\mathcal{G}_{n}\right)\right)=\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{G}\right)}\left(\eta^{\left(0\right)}\right)\right)\leq\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{G}\right)}\right)+1\leq n-\lfloor n/2\rfloor+2,$
+where the final inequality follows from (24).
+Moreover, we know from Theorem 2 that the values in $\mathcal{H}_{n}^{\prime}$ are algebraically independent, implying $\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{H}^{\prime}\right)}\right)=|\mathcal{H}_{n}^{\prime}|=n+1$ and therefore
+$n+1\leq n-\lfloor n/2\rfloor+2+\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{D}^{\prime}\right)}\right)$
+$\Longleftrightarrow\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{D}^{\prime}\right)}\right)\geq\lfloor n/2\rfloor-1.$
+Since $\underset{n\rightarrow\infty}{\lim}\left(\lfloor n/2\rfloor-1\right)=\infty$, it follows that the sequence $\left\{\delta^{\left(n\right)}\right\}_{n\geq 1}$ must contain an infinite number of transcendental terms.
+(iv) The density of the set of transcendental $\delta^{\left(n\right)}$ evaluated at $N\in\mathbb{Z}_{\geq 2}$ satisfies
+$\dfrac{\#\left\{n\leq N:\delta^{\left(n\right)}\textrm{ is transcendental}\right\}}{N}\geq\frac{\mathrm{trdeg}\left(\mathcal{F}_{N}^{\left(\mathcal{D}^{\prime}\right)}\right)}{N}\geq\dfrac{\lfloor N/2\rfloor-1}{N},$
+with asymptotic lower bound
+$\underset{N\rightarrow\infty}{\liminf}\left(\dfrac{\lfloor N/2\rfloor-1}{N}\right)=\dfrac{1}{2}$
+as $N\rightarrow\infty$. ∎
 4. Non-Alternating Analogues
@@
 $\gamma+\dfrac{\widetilde{\delta}}{e}=-{\displaystyle\sum_{k=1}^{\infty}\dfrac{1}{k\cdot k!}}=\textrm{Ein}\left(-1\right),$ (29)
-which is shown to be transcendental by Theorem 4 of Section 4.3 below.
+which is shown to be transcendental by Theorem 4 of Section 4.3.
 4.2. Asymptotic Behavior
@@
 However, the (consistently negative) $\widetilde{\delta}^{\left(n\right)}$ values are quite different in magnitude from the $\delta^{\left(n\right)}$ of Table 1 because they, like the $\widetilde{\eta}^{\left(n\right)}$, grow approximately factorially as well.
-Table 2. Values of $\widetilde{\delta}^{\left(n\right)}$ and $\widetilde{\eta}^{\left(n\right)}$ for $n\in\left\{0,1,\ldots,15\right\}$
-$\boldsymbol{n}$ $\boldsymbol{\widetilde{\delta}^{\left(n\right)}}$ $\boldsymbol{\widetilde{\eta}^{\left(n\right)}}$
-0 7.3890560989-7.3890560989 1.7182818285-1.7182818285
-11 5.1514643230-5.1514643230 1.3179021515-1.3179021515
-22 11.6100810693-11.6100810693 2.2929981451-2.2929981451
-33 32.2422478187-32.2422478187 6.4163856532-6.4163856532
-44 131.4700021171-131.4700021171 24.8036368256-24.8036368256
-55 651.8492520020-651.8492520020 121.9625302861-121.9625302861
-66 3,916.6763381264-3,916.6763381264 725.7973399920-725.7973399920
-77 27,400.1009939266-27,400.1009939266 5,060.0849690569-5,060.0849690569
-88 219,211.5495238585-219,211.5495238585 40,399.8007638273-40,399.8007638273
-99 1,972,830.5386794810-1,972,830.5386794810 363,237.5069807080-363,237.5069807080
-$10$ 19,728,269.2785592814-19,728,269.2785592814 3,630,582.2647192273-3,630,582.2647192273
-$11$ 217,010,407.4543390336-217,010,407.4543390336 39,926,583.2712579089-39,926,583.2712579089
-$12$ 2,604,123,546.6077724787-2,604,123,546.6077724787 479,060,223.3022788290-479,060,223.3022788290
-$13$ 33,853,598,434.1585762781-33,853,598,434.1585762781 6,227,401,522.0546076015-6,227,401,522.0546076015
-$14$ 473,950,346,279.9399036234-473,950,346,279.9399036234 87,180,954,721.7674533138-87,180,954,721.7674533138
-$15$ 7,109,255,026,416.1913290094-7,109,255,026,416.1913290094 1,307,694,336,767.4617097988-1,307,694,336,767.4617097988
 The following proposition quantifies the asymptotic behavior of both sequences.
@@
 See the Appendix. ∎
+Table 2. Values of $\widetilde{\delta}^{\left(n\right)}$ and $\widetilde{\eta}^{\left(n\right)}$ for $n\in\left\{0,1,\ldots,15\right\}$
+$\boldsymbol{n}$ $\boldsymbol{\widetilde{\delta}^{\left(n\right)}}$ $\boldsymbol{\widetilde{\eta}^{\left(n\right)}}$
+0 $7.3890560989\ldots$ $1.7182818285\ldots$
+11 $5.1514643230\ldots$ $1.3179021515\ldots$
+22 $11.6100810693\ldots$ $2.2929981451\ldots$
+33 $32.2422478187\ldots$ $6.4163856532\ldots$
+44 $131.4700021171\ldots$ $24.8036368256\ldots$
+55 $651.8492520020\ldots$ $121.9625302861\ldots$
+66 $3,916.6763381264\ldots$ $725.7973399920\ldots$
+77 $27,400.1009939266\ldots$ $5,060.0849690569\ldots$
+88 $219,211.5495238585\ldots$ $40,399.8007638273\ldots$
+99 $1,972,830.5386794810\ldots$ $363,237.5069807080\ldots$
+$10$ $19,728,269.2785592814\ldots$ $3,630,582.2647192273\ldots$
+$11$ $217,010,407.4543390336\ldots$ $39,926,583.2712579089\ldots$
+$12$ $2,604,123,546.6077724787\ldots$ $479,060,223.3022788290\ldots$
+$13$ $33,853,598,434.1585762781\ldots$ $6,227,401,522.0546076015\ldots$
+$14$ $473,950,346,279.9399036234\ldots$ $87,180,954,721.7674533138\ldots$
+$15$ $7,109,255,026,416.1913290094\ldots$ $1,307,694,336,767.4617097988\ldots$
 4.3. Transcendence Results
@@
 Theorem 5.
-For the sequence $\left\{\widetilde{\delta}^{\left(n\right)}/e\right\}_{n\geq 1}$:
-(i) $\widetilde{\delta}^{\left(n\right)}/e$ is transcendental infinitely often; and
-(ii) the density of the set of transcendental $\widetilde{\delta}^{\left(n\right)}/e$ (i.e., $\left\{n\leq N:\widetilde{\delta}^{\left(n\right)}/e\textrm{ is transcendental}\right\}/N$) has asymptotic lower bound $1/2$ as $N\rightarrow\infty$.
+For the sequences $\left\{\widetilde{\delta}^{\left(n\right)}/e\right\}_{n\geq 1}$ and $\left\{\widetilde{\delta}^{\left(n\right)}\right\}_{n\geq 1}$:
+(i) $\widetilde{\delta}^{\left(n\right)}/e$ is transcendental infinitely often;
+(ii) the density of the set of transcendental $\widetilde{\delta}^{\left(n\right)}/e$ (i.e., $\left\{n\leq N:\widetilde{\delta}^{\left(n\right)}/e\textrm{ is transcendental}\right\}/N$) has asymptotic lower bound $1/2$ as $N\rightarrow\infty$;
+(iii) $\widetilde{\delta}^{\left(n\right)}$ is transcendental infinitely often; and
+(iv) the density of the set of transcendental $\widetilde{\delta}^{\left(n\right)}$ (i.e., $\left\{n\leq N:\widetilde{\delta}^{\left(n\right)}\textrm{ is transcendental}\right\}/N$) has asymptotic lower bound $1/2$ as $N\rightarrow\infty$.
 ?proofname?.
-The proofs of both parts of this theorem are entirely analogous to those of the corresponding parts of Theorem 3. ∎
+The proofs of all parts of this theorem are entirely analogous to those of the corresponding parts of Theorem 3. ∎
 5. Conclusion
@@
 After characterizing the basic asymptotic behavior of these sequences, we provided the following (unconditional) results:
-$\bullet\>$ at least one element of the pair $\left\{\gamma,\gamma^{\left(2\right)}\right\}$ and at least two elements of each set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$, for $n\in\left\{2,3,\ldots,26\right\}$, are transcendental, with analogous implications for the sequence $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$;
+$\bullet\>$ the $\gamma^{\left(n\right)}$ are transcendental infinitely often and at least two elements of each set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$, for $n\in\left\{2,3,\ldots,26\right\}$, are transcendental (with analogous implications for the sequence $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$);
 $\bullet\>$ the $\eta^{\left(n\right)}$ are algebraically independent, and therefore transcendental, for all $n\geq 0$;
 $\bullet\>$ at least one element of each pair, $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e\right\}$ and $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}\right\}$, and at least two elements of the triple $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e,\delta^{\left(n\right)}\right\}$, are transcendental for all $n\geq 1$; and
-$\bullet\>$ the $\delta^{\left(n\right)}/e$ are transcendental infinitely often, with the density of the set of transcendental terms having asymptotic lower bound $1/2$.
+$\bullet\>$ both the $\delta^{\left(n\right)}/e$ and $\delta^{\left(n\right)}$ are transcendental infinitely often, with the densities of both sets of transcendental terms having asymptotic lower bound $1/2$.
 Subsequently, for the generalized non-alternating Eta and Euler-Gompertz constants, denoted by $\widetilde{\eta}^{\left(n\right)}$ and $\widetilde{\delta}^{\left(n\right)}$, respectively, we found:
@@
 $\bullet\>$ at least one element of each pair, $\left\{\gamma^{\left(n\right)},\widetilde{\delta}^{\left(n\right)}/e\right\}$ and $\left\{\gamma^{\left(n\right)},\widetilde{\delta}^{\left(n\right)}\right\}$, and at least two elements of the triple $\left\{\gamma^{\left(n\right)},\widetilde{\delta}^{\left(n\right)}/e,\widetilde{\delta}^{\left(n\right)}\right\}$, are transcendental for all $n\geq 1$; and
-$\bullet\>$ the $\widetilde{\delta}^{\left(n\right)}/e$ are transcendental infinitely often, with the density of the set of transcendental terms having asymptotic lower bound $1/2$.
+$\bullet\>$ both the $\widetilde{\delta}^{\left(n\right)}/e$ and $\widetilde{\delta}^{\left(n\right)}$ are transcendental infinitely often, with the densities of both sets of transcendental terms having asymptotic lower bound $1/2$.
 Although our primary focus has been to investigate properties of the sequences $\gamma^{\left(n\right)}$, $\delta^{\left(n\right)}$, $\eta^{\left(n\right)}$, $\widetilde{\delta}^{\left(n\right)}$, and $\widetilde{\eta}^{\left(n\right)}$, it is important to note that further insights may be obtained by decomposing the sequence of $E$-functions in (23) by the method of multisections.
@@
 $H_{n}\left(z\right)=-n!{\displaystyle\sum_{k=2,4,\ldots}\dfrac{z^{k}}{k^{n}\cdot k!}},$
-each of which is itself an $E$-function, and these new functions arranged into a system of linear differential equations with coefficients in $\mathbb{Q}\left(z\right)$ and a unique finite singularity at $z=0$.
+each of which is itself an $E$-function, and these new functions can be arranged into a system of linear differential equations with coefficients in $\mathbb{Q}\left(z\right)$ and a unique finite singularity at $z=0$.
 By employing the Kolchin-Ostrowski differential-field framework (see Srinivasan [18]), we then can show that, for any $m\in\mathbb{Z}_{\geq 1}$, the functions $G_{1}\left(z\right),H_{1}\left(z\right),G_{2}\left(z\right),H_{2}\left(z\right),\ldots,G_{m}\left(z\right),H_{m}\left(z\right)$ are algebraically independent over $\mathbb{C}\left(z,G_{0}\left(z\right),H_{0}\left(z\right)\right)$.
@@
   • [3] Mahler, K., 1968, “Applications of a Theorem by A. B. Shidlovski”, Proceedings of the Royal Society A, 305, 1481, 149-173.
-  • [4] Rivoal, T., 2012, “On the Arithmetic Nature of the Values of the Gamma Function, Euler’s Constant, and Gompertz’s Constant”, Michigan Mathematics Journal, 61, 2, 239-254.
+  • [4] Rivoal, T., 2012, “On the Arithmetic Nature of the Values of the Gamma Function, Euler’s Constant, and Gompertz’s Constant”, Michigan Mathematical Journal, 61, 2, 239-254.
   • [5] Lagarias, J. C., 2013, “Euler’s Constant: Euler’s Work and Modern Developments”, Bulletin of the American Mathematical Society, 50, 4, 527-628.
   • [6] Hardy, G. H., 1949, Divergent Series, Clarendon Press, Oxford, UK.
-  • [7] Powers, M. R., 2025, “Conjunctions of Three ’Euler Constants’ in Poisson-Related Expressions”, arXiv:2503.02054.
+  • [7] Powers, M. R., 2025, “Conjunctions of Three ’Euler Constants’ in Poisson-Related Expressions”, arXiv:2503.02054v4.
   • [8] Pinkus, A., 2010, Totally Positive Matrices, Cambridge University Press, Cambridge, UK.
@@
   • [15] Beukers, F., 2006, “A Refined Version of the Siegel-Shidlovskii Theorem”, Annals of Mathematics, 163, 369-379.
-  • [16] Powers, M. R., 2025, “On the Uniqueness of Ein(1) among Linear Combinations of the Euler-Mascheroni and Euler-Gompertz Constants”, arXiv:2510.00315.
-  • [17] Berndt, B. C. and Rankin, R. A., 1995, Ramanujan: Letters and Commentary, American Mathematical Society.
-  • [18] Srinivasan, V. R., 2008, “Extensions by Antiderivatives, Exponentials of Integrals, and by Iterated Logarithms”, arXiv:0811.3004.
+  • [16] Powers, M. R., 2025, “On the Uniqueness of Ein(1) among Linear Combinations of the Euler-Mascheroni and Euler-Gompertz Constants”, arXiv:2510.00315v1.
+  • [17] Berndt, B. C. and Rankin, R. A., 1995, Ramanujan: Letters and Commentary, American Mathematical Society, Providence, RI, USA.
+  • [18] Srinivasan, V. R., 2008, “Extensions by Antiderivatives, Exponentials of Integrals, and by Iterated Logarithms”, arXiv:0811.3004v1.
   • [19]